A Complete Bibliography of Publications in the Journal of Statistical Planning and Inference:

2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

06 April 2024 Version 1.19

Title word cross-reference

2 [HLYZ20]. 2^{K} [Gro20, DHS21]. 2^{m} [KH20]. 2ℓ [KH20]. 2r [HLYZ20]. A [LZ20a, PS21]. α [FHM22, Has21]. c [PS21]. D [GZ20, HH21, WY22]. ℓ_{1} [DT22]. ℓ_{4} [LX23]. I [Hai22]. K [MLM21, SST20, TLG24]. L_{0} [LPK^{+}21]. L_{2} [ZYL20]. M [CM23, KV21]. p [CS20, Hil24]. ψ [KW24]. s∗ [LMW21]. t [BS22, MAC22]. U [FHBD20, HLP23, Kim20, KS22, SM23].

-and [PS21]. -concave [LMW21]. -distance [ZYL20]. -divergence [Has21]. -estimation [CM23]. -factorial [DHS21]. -level [HLYZ20]. -norm [LX23]. -optimal [GZ20, Hai22, HH21, LZ20a, PS21, WY22]. -sample [TLG24]. -shrinkage [BS22]. -sign [MLM21]. -stable [FHM22]. -statistic [FHBD20, SM23]. -statistics [HLP23, Kim20, KS22]. -type [KV21]. -value [Hil24]. -values [CS20]. -variations [SST20]. -walk [MAC22]. -weakly [KW24].

2020 [Ano20k, Ano20j]. 2021 [Ano21k, Ano21j, Ano21l, Ano21i, Ano21g]. 2022 [Ano22j, Ano22h, Ano22k, Ano22i, Ano22g, Ano22l]. 2023 [Ano23i, Ano23j, Ano23k, Ano23l, Ano23h, Ano23g]. 2024 [Ano24g, Ano24h, Ano24i, Ano24k].

aberration [ZZ20]. aberrations [BS21a]. ability [FHS23]. accelerated [QW23, SLX20, TLT22]. actions [SMB20]. active [LFW22]. activity [BNF^{+}23, Mie21]. adaptation [FMT21]. Adapting [NSC21]. Adaptive [CCHR22, GS20, KU22, AG21, DT24, DZ24, GZZ20, KY24, Lan23, LC20, MCJ23, SW22, WVJ^{+}20, WLFL24, ZCQC21]. Adaptively [WZSQ24]. addition [Tsa22, WST21, WM22]. additional [Wal23]. additive [GMFY20, GC21, LL23, SA20, WMZ20, WLFL24]. adequacy [MR21, SDSZ21]. Adjusted [AS22, MCJ23, ZY22]. adjustment [MR21]. admissibility [MS23, Su22]. admixture [MPSV22]. against [KKSA22, PSS20]. age [AFMY20, BKM21]. aggregation [SZ22]. aids [LFR20]. algorithm [GPP23, LCYZ21, OYF20, YCL21]. algorithmic [KVH^{+}23]. algorithms [CIO24]. allow [TLG24]. allowing [WZC^{+}21]. almost [LMMJ21]. Alternative [KCD23, Hil24]. alternatives [PSS20]. analyses [EGC20]. Analysis [BdFL20, Tan22, Wal23, ASN22, AP21, Cav22, DJPZ22, DCDS21, FLL^{+}24, FHL24, HLLW24, KU24, LW20, LFR20, MG23, SSC20, SCCF24, SLW24, VdC21, WZSQ24, YHDB22, ZC20b]. Analytical [TLT22]. Analyzing [GM21]. Andersen [CSL20]. Anderson [Zha21]. Anisotropic [Dus21, AG21]. any [Hua21]. Application [DT22, CM23, GPG20a, GPG20b, LAO22, Pru22, WZL^{+}22]. applications [AGNUC23, BKP20, BB21, BNF^{+}23, DT24, GZ23, JHH21, LYZ21, LPK^{+}21, Oui21, ZZC24]. applied [CS20]. approach [AFMY20, DS20a, GC21, Gór24, JT23, KVH^{+}23, LLBR22, LZW^{+}24, MS21a, NMP21, NBD21, SW22, TLT22, ZZGZ21]. Approximate [Hai22, LZ20b, PFM21, RSV20]. Approximating [BFVT22]. approximation [XMX21]. Archimedean [HANB20]. area [TJ20, WM20]. areas [WK23]. ARMA [SML22]. armed [AHY21]. arrays [CT22a, Hua21, LLY21, LMMJ21, WY22]. assemblies [KH20]. assets [WZL^{+}22]. assignment [GPG20a, GPG20b, ZCQC21]. assisted [GR24, SD21, WD23]. association [DJPZ22, WT24]. assumption [Pil20, TL24]. aster [EGC20]. asymmetric [HYD24]. asymmetry [FZ24]. Asymptotic [BOAB20, CG22, DSK21, KS22, KvW23, BL20, FHBD20, Kim20, KCD23, YH23]. Asymptotically [HB21, Rya22, BBCS24]. asymptotically-efficient [BBCS24]. Asymptotics [BKP20, CB20, LW22, MB20, MLM21]. asynchronous [SCCF24]. attribute [LTW21]. augmented [WY22]. autocovariance [LLBR22]. autoregression [DK21]. autoregressive [BOAB20, DR24, GS24, GL20, BBCS24, LY20, SWS24, YWYL20, ZWYX20]. auxiliary [FLL^{+}24, KV21]. average [DR24, EMM21, WXT^{+}21, ZZ21]. averaged [KWW20]. averaging [BJZ22, HCZG24, ZZWZ22, ZZ24]. balance [ZY22]. balanced [DLZ23, GM21, JT23]. ball [Hai22]. ballots [GPG20a, GPG20b]. bandit [AHY21]. bands [TS22]. Bandwidth [CM20]. based [AGNUC23, AM23, BI22, CWY20, DT22, FDGD24, FC20, GD20, GKO22, HCX23, Jeo23, KU21, KU22, Kim20, KS22, KRX21, LCKB24, LX23, LJ20, MU21, SWS24, SM23, WWZZ20, WXT^{+}21, XC20, YH23, Yan20, ZDW^{+}21, Zha21, ZLW^{+}24, ZZ21, ZZ24]. baseline [LLT22]. bases [FC20]. basis [Dus21]. Bayes [IK21, JX22, YK23]. Bayesian [AG20, AGB20, AGNUC23, AM23, BG22a, BG22b, BFVT22, DCDS21, Gór24, GR24, HK22, HH24, HANB20, LAYT22, LG21, LGM23, MB20, MK21, MG23, Ove22, RD21, RSV20, VdC21, XMX21, YHDB22, ZH23, ZYG22]. behaviour [FHBD20]. Behrens [ZZGZ21]. Benjamini [CS20]. bent [WZL^{+}22]. Berman [ASN22]. Bernoulli [FS23, LPL21]. best [ZCQC21]. beta [AM23]. beta-Stacy [AM23]. between [FC20, KZ21]. Beyond [ACCY21, FZ24]. Bi [LMW21, KvW23]. Bi- [LMW21]. bi-section [KvW23]. bias [GPG20a, GPG20b, KZ21, NMP21]. biased [Jaz24, LCKB24]. bidimensional [BOAB20]. bifurcating [Pen24]. big [CWY20]. binary [CBB24, HHL20, LFW22, WF20]. binomial [HH24, LJ20, SWS24, ZWL22]. Birnbaum [DGCL21, Kak20]. Bivariate [YWYL20, BKP20, DMK22]. block [BCF^{+}24, DLZ23, Gro20, GS20, SS20, UB22, YFS22]. blocked [Gro20, TH22]. BMM [BOAB20]. BMM-estimator [BOAB20]. Board [Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano24a, Ano24b, Ano24c, Ano24d, Ano24e]. bootstrap [AM23, HLP23, NBD21]. both [YYYW21]. bound [KH20, Lop20]. bounded [YFS22]. Box [CQY22]. breaking [RD21]. breathing [BNF^{+}23]. Bregman [DCDS21]. brief [Don21]. Brown [MS23]. buffered [LY20]. business [BB21].

Calculating [MP21]. calculation [LFLL21]. calculus [SST20]. Calibrating [LWL^{+}24]. calibration [CG22, JT23]. canonical [LZ20a, Sal21, ZC20b]. card [KP24]. cards [KP24]. Carlo [MM20]. CARMA [FHM22]. case [BS21b, GC20, MPSV22, QW23]. case-cohort [QW23]. categorical [DCS20, HS24b, Jeo22]. causal [GR24, LGM23, WXT^{+}21]. cautionary [SSdW22]. censored [AM23, Che24, LC20, MCJ23, QW23, SLX20]. censoring [HLDB22, KN24, ZFZ20]. centered [KWW20]. Central [LYZ21, Kak20]. certain [Has21]. chains [Pen24]. Change [WD23, KS22, LG21, LCYZ21, LZ22, ZH23, ZYLY22]. Change-detection-assisted [WD23]. change-point [LCYZ21, LZ22]. changed [GL20]. changes [GS24, JK20]. characteristic [WM20]. characteristics [BFVT22]. Checking [SDSZ21, YYW23]. chi [HK22]. chi-squared [HK22]. Choice [GKO22, DS20a, Gro20]. circulant [LMMJ21]. Circular [KZZ23, FDGD24, MK21]. class [CCHR22, KY24, Pro20, SA20]. classes [AG21]. classical [MB20]. classified [NSC21, SLJ20]. classifiers [Wal23]. clear [HLYZ20]. clinical [GZZ20, HI22, YYYW21]. closed [YK23]. cluster [LW22]. clustered [DSK21, HC23]. coding [DCS20]. coefficient [FTHL21, HDLF20, LCP22, YWYL20]. coefficients [GKO22, GHRS20, XZ21]. cohort [AFMY20, QW23]. cokriging [DMK22]. collinearity [BS21c, KKH24]. Column [LLY21, KVH^{+}23]. Column-orthogonal [LLY21]. combinatorial [Fro22]. Combining [EGC20]. communities [KvW23]. commutative [BCF^{+}24]. Comparing [ZCQC21, BKM21, HYD24]. comparison [GD20, Gro20, LCKB24, LCZ24, MPSV22]. comparisons [BL22, NLL20, ZCQC21]. compatible [Llo22]. competing [AKE22, WZC^{+}21]. complement [JS22]. complementary [ZZ20]. completely [BdFL20, Buo20]. complex [LWL^{+}24, NS24]. component [AP21, BB24, CCHR22, GPP23, Hua21, LL23]. components [BFR21, Hua21, MPSV22, PdOA22, RX20, WL24]. Composite [PX24, BL20, ZLZ20]. compound [MU21]. compromise [LLT22]. computationally [DZ24]. computer [MM20]. concave [BB24, LMW21]. concentration [AG20]. concentric [ASN22]. concordance [Che24]. condition [LYH20]. conditional [CBL20, FC20, FZ24, HCX23, YM24, ZZ21]. conditionally [dSCBP23]. conditions [BBS21]. Confidence [KC24, DJPZ22, GAC23, HS24b, IK21, Jeo23, KWW20, Llo22, LP21, PFM21, TS22]. confounding [KLCZ23, SZ23]. connected [KLR23]. connectivity [BKP20]. Consistency [DS20b, Pil20, CWS20, NLL20]. consistent [CFM23]. constrained [XMO21]. constraint [LFLL21]. constraints [YZW23, ZZWZ22]. Constructing [ZZ20, KLCZ23]. Construction [HS24a, HWS24, Hua21, LYCL24, WYL21, ZXS23, AGB20, BDL^{+}23, KVH^{+}23, LLT22, QLZ23, WM24]. Constructions [UB22]. content [TL24]. contingency [SSdW22]. continuous [LW20]. Contraction [WG20, ZYG22, Jeo22]. contrasts [Ros20]. control [Cla21, HB21, XLZ20, YYYW21]. controlling [NSC21, NS24]. controls [GS20]. Convergence [GZ23, Lop20]. convolutional [Wal23]. convolved [SSC20]. copula [GKO22, HANB20]. copula-based [GKO22]. copulas [KY24, ZWL22]. Corrected [IK21]. Correcting [SSdW22]. correction [NMP21]. correlated [DLZ23, GHRS20, SKS21]. correlation [HS24b, LX23, WF20, Yin20, YW20, ZC20b]. correlations [NKST21]. correntropy [YJL24]. corresponding [KZ21]. Corrigendum [GPG20a]. count [BAA24, SS20, WWZZ20]. counts [SSC20, ZWYX20]. covariance [BM21, CSP21, LYZ21, SBP22, Su22, WL20, YZZ22, YCL21, YK23, ZHB20]. Covariate [MCJ23, BS21c, ZYLY22]. Covariate-adjusted [MCJ23]. covariates [AG20, AHY21, Chu23, GMFY20, GR24, LPL21, Ros20, SCCF24, XDZL20, YW20, ZSS23]. COVID19 [HLDB22]. Cox [CQY22, Jaz24, LB22, QLXL24]. Cressie [RW23]. criterion [KKSA22, OYF20, YJL24]. cross [GM21]. cross-balanced [GM21]. crossing [LW20]. crossover [AHB21]. cumulative [AKE22]. curable [WZC^{+}21]. cure [WZC^{+}21]. curve [RKvS24, WM20]. curves [HYD24, WK23].

D4R [MWL24]. Darling [Zha21]. Darmois [CLPX23]. data [AM23, CSL20, CG22, Che24, CWY20, Chu23, CF20, DCDS21, DZ24, DSK21, FDGD24, FHL24, GC20, GD20, HC23, HH24, HH21, JT23, Jaz24, Kak20, KU22, KY24, KRX21, LYH20, LG21, LPK^{+}21, LWL^{+}24, LY21, LYH22, EFR20, QLXL24, QW23, Sae22, SS20, SSC20, SA20, SCCF24, SLX20, Tan22, TS22, WZL20, WWZZ20, WD23, XMO21, ZZGZ21]. data-adaptive [DZ24, KY24]. datasets [YYW23]. December [Ano20i, Ano21k, Ano22j, Ano23i]. decomposition [DLZ23]. deconvolution [Dus21]. Deep [KW24, KLR23, LWL^{+}24]. default [MSFT23]. definite [RKvS24]. degradation [TLT22]. degrees [Gór24, MP21, YFS22]. delay [BP21, KS22]. denoising [BS22]. densities [Ste22]. Density [TTML21, AG21, BB24, CKSW21, CCHR22, CMZ20, CM23, HK22, Kak20, MS21b, Pen24, RD21, ZHLT20]. dependence [dSCBP23, BNF^{+}23, FZ24, GS20, HCX23, KZ21, LP21]. dependency [NBD21]. dependent [Chu23, KW24, Mie21, MM20, MB24]. depth [MLM21]. design [BS21c, FHS20, FHS23, GM21, HC23, LTZF22, Ove22, PL20, PS21, Pru22, WVJ^{+}20, ZCQC21]. designing [TLT22]. Designs [BCF^{+}24, BFVT22, AHB21, BFR21, CT22b, DMK22, DS20b, DHS21, DLZ23, FHS20, GZ20, GHRS20, Hai22, HLYZ20, HY20, HS24a, HYD24, HH21, HWS24, HHL20, KKSA22, KVH^{+}23, KZZ23, KLCZ23, KH20, Lan23, LZ20a, LLT22, LYCL24, LC20, LMMJ21, MCJ23, NZ22, PW21, QLZ23, Ros20, Sab21, SBP22, SS20, SKS21, SZ23, TL24, Tsa22, TH22, WYL21, WY22, WCL22, WM22, WM24, XS21, YYYW21, YZZ22, ZLZ20, ZZ20, ZYL20, ZXS23]. Detecting [LTZF22]. detection [CF20, LG21, WD23, ZH23]. deteriorating [SMB20]. Deterministic [QLZ23]. deviations [Fro22, Pro20]. diagnostic [KZ21]. diallel [BCF^{+}24]. difference [SZ21, XC20]. differential [BP21, Cla21, MB20]. diffusion [AG21, KU22, KU24, MU21, TN23]. dimension [Don21, DSP23, MWL24, SD21]. dimensional [CSP21, CG22, DJPZ22, FS23, FHL24, GC21, HCX23, HS24a, JHH21, Kim20, LX23, LWL^{+}24, LZW^{+}24, LZ22, LL23, WMZ20, WL23, WZSQ24, WC24, WL20, XLZ20, XZ21, YH23, ZHB20, ZZGZ21]. dimensions [BM21, Su22]. direct [YCL21]. directed [BFVT22]. Directional [WZL20, BG22b]. Dirichlet [LB22]. discovery [NSC21, NS24]. Discrete [DS20a, BM20, Cla21, KU21, KU24, TN23]. discrete-valued [BM20]. discretely [Kal23]. discreteness [TPZ22]. disease [YYYW21]. disorder [Buo20]. dispersion [HH21, TLT22]. dispersive [Sal21]. distance [GK22, HS24a, ZYL20]. Distributed [FLL^{+}24]. distribution [AVVW24, AGNUC23, DGCL21, DGCL22, EMM21, GC20, JX22, KZA23, Oui21, Sae21, TTML21, YH23]. distributional [AHZ20, Yan20]. distributions [BS21b, CF24, FS23, GAC23, Khm21, KC24, LMW21, LAO22, LB22, MB20, MSFT23]. divergence [DCDS21, Has21, XC20]. domain [BL20, MR21, SW22, WH20, XMO21]. dominance [Gór24]. dose [CIO24, LFR20]. double [GL20, HLDB22]. double-censoring [HLDB22]. Doubly [MWL24, ZHLT20, LGM23]. drift [Mie21]. driven [CLT23]. due [ZYLY22]. dynamic [LYH22]. dynamical [dSCBP23].

E-variables [TLG24]. early [CIO24, KS22]. Editorial [Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano24a, Ano24b, Ano24c, Ano24d, Ano24e]. effect [YW20]. effects [AHB21, AHZ20, AFMY20, BFR21, yCjL23, FMT21, FHS20, Gro20, HLYZ20, KZ21, KZZ23, MB20, SS20, WXT^{+}21, ZZ21]. Efficiency [Lan23, CG22, SMB20, WK23]. Efficient [BS21b, HLDB22, LW20, Nar21, YCL21, DZ24, BBCS24, JT23, LMMJ21, MLM21, WM24]. Efron [BFMS21]. eigenspace [CKSW21]. eigenvalues [JHH21]. elastic [LAYT22, Pil20]. election [GPG20a, GPG20b]. elliptical [DGCL21, DGCL22]. Empirical [WL23, FHL24, GK22, HS24b, HLDB22, IK21, KY24, SZ21, ZLW^{+}24]. empirically [TJ20]. endpoints [LW20]. energy [MSR21]. ensembles [Lop20]. entropy [HI22]. envelope [EGC20, ZC20a]. eQTL [FLL^{+}24]. equation [SST20]. equations [BP21, Cla21, DSK21, MB20]. Equivalence [Pru22]. equivariant [BM21]. ergodic [AG21, KU22]. error [LW22, TJ20, TTML21]. errors [DLZ23, FMT21, GK22, HYD24, HCZG24, SSdW22, SKS21, ZSS23]. errors-in-variables [GK22]. estimate [BI22, LLBR22, SMB20]. estimates [GZ23]. Estimating [LZ22, Lop20, MU21, NBD21, BFR21, DSK21, FHS23]. Estimation [BFMS21, EMM21, FTHL21, GMFY20, Jaz24, KLR23, Mie21, TJ20, XDZL20, AKE22, ASN22, AG21, BdFL20, BB24, BS21b, CKSW21, CCHR22, CM23, CFM23, Cla21, DJPZ22, DHS21, FHM22, FS23, GZZ20, GK21, GL20, HK22, BBCS24, HH21, KU21, Kal23, KV21, LPL21, LYH20, LCYZ21, LZW^{+}24, LZ20b, LGM23, MRS21, MS21b, MT23, MMD20, MB24, MSFT23, Nar21, Pen24, SML22, SST20, SD21, SLX20, TPZ22, TTML21, WST21, WZC^{+}21, WF20, WY20, WXT^{+}21, XMO21, YX23, YCL21, ZHLT20, ZYG22, ZZ21]. estimator [BJZ22, BM21, BOAB20, DT24, CG22, CLT23, CM20, DS20b, FHBD20, KWW20, Kak20, EFR20, NMP21]. estimators [BL20, BKP20, CWS20, GK22, GD20, GPP23, Jeo23, KZ21, RKvS24, XMO21, YK23]. EV [CWS20]. Evaluation [CIO24]. event [CSL20, LW20, SA20]. evidence [PdOA22]. evolving [DSK21]. Exact [BL22, Llo22, Jeo23, TLG24]. Exchangeable [FS23, ZH23]. existence [WY22]. expansion [YH23, ZXS23]. expansive [LYCL24]. expected [MB24]. expectile [AGNUC23, BJZ22, SD21]. expectile-assisted [SD21]. expectile-based [AGNUC23]. experimental [BS21c, DS20b, PW21, Ros20]. experiments [BCF^{+}24, DS20a, Gro20, JS22, KKSA22, LMMJ21, MM20, Ove22]. exponential [KN24, LZ20b, MP23, MRS21, TLT22]. extended [GS24]. extending [YM24]. extensions [yCjL23]. extremal [KOV22]. factor [KU24, LZ22, WZSQ24]. factorial [DHS21, HLYZ20, KH20, Sab21]. factors [BFR21, GZ20, JX22, SZ23, ZZ20]. failed [KKSA22]. failure [BKM21, EMM21, QW23]. false [NSC21, NS24]. families [MP23]. family [BP21, CLPX23, MRS21]. Fast [BBCS24, WM24, OYF20]. faster [ZLHZ22]. Fay [AS22, IK21]. Fay-Herriot [AS22]. FDR [GS20, HB21]. Feature [Che24, LL23, WMZ20, DZ24, GC21]. Fields [LZ20b]. files [GR24]. filling [CT22b, WYL21]. filtering [Rez22]. financial [MB24, WZL^{+}22]. finding [CIO24]. Finite [KWW20, GAC23, Pil20, ZWYX20]. finite-range [ZWYX20]. First [SWS24, YWYL20]. First-Order [SWS24, YWYL20]. Fisher [JHH21, Ove22, ZZGZ21]. fit [Cav22, Kim20, PSS20, XC20, ZSS23]. Fitting [LY21, YW20]. Fixed [BL20, MSFT23, TN23]. Fixed-domain [BL20]. flats [WM24]. Flexible [ZWL22, Tsa22]. FMANOVA [ZZC24]. fMRI [LMMJ21, SKS21]. forms [YK23]. formula [TS22]. fractal [BKP20]. fractional [HLYZ20, BBCS24, Kal23, KH20, NMP21, Nar21, Sab21, SST20]. fractionally [DR24]. framework [BL22, HH24, KKH24, Sab21]. free [DZ24, WM20, Yin20]. freedom [MP21]. frequencies [FHM22]. frequency [KU22, LL23, MR21]. Frequentist [ZZWZ22, PdOA22]. fully [KLR23]. function [EMM21, GC20, KCH21, KLR23, LLBR22, WST21, WZSQ24, XLZ20]. functional [AG20, BNF^{+}23, CF20, FTHL21, GZ23, HDLF20, HXZZ20, KRX21, LG21, LLZL21, LYH22, PSS20, RX20, SDSZ21, SW22, SLW24, TS22, WZL20, XDZL20, XLZL22, ZSS23, ZLHZ22]. functional/longitudinal [LYH22]. functions [AKE22, BKM21, LL23, Sae21, Su22, WWZZ20]. Further [NS24]. FWER [GS20].

gain [Ove22]. gamma [LB22, SMB20, SS20]. Gaussian [AGB20, BL20, CLT23, CMMD21, CBL20, MS21a, MS23, SSC20, TS22]. GED [Rez22]. General [SZ23, AM23, Baz23, DS20b, DK21, HLLW24, KLCZ23, LYZ21, PSS20, SBP22, SA20, YH23, ZCQC21]. Generalization [ZDW^{+}21]. generalize [DCDS21]. Generalized [AP21, QW23, SST20, YK23, ZYLY22, GAC23, JHH21, LAYT22, LPK^{+}21, OYF20, SW22, XS21, ZY22]. generate [CT22b]. generated [Chu23, GMFY20]. Generating [Tsa22]. generators [Rya22]. Generic [TLG24]. genome [DJPZ22, WT24]. genome-wide [DJPZ22, WT24]. geometric [Sab21]. Gibbs [BM22]. Gill [CSL20]. Gini [HS24b]. given [BP21]. global [LNH^{+}23, ZZC24]. Goodness [Cav22, ZSS23, Kim20, XC20]. Goodness-of-fit [Cav22, ZSS23, Kim20, XC20]. Graph [BS22, DLZ23, GD20, UB22]. Graph-based [GD20]. graphical [MG23]. greedy [LCYZ21]. Group [CSL20, LCYZ21, LC20, CB20, DCS20, Jeo22, NZ22, SW22, ZHLT20]. Group-sequential [LC20]. grouping [NS24]. guessing [KP24].

Hadamard [HWS24]. half [BCF^{+}24]. half-diallel [BCF^{+}24]. having [MB20]. Hawkes [DT24, KCD23]. hazard [Tan22]. hazards [QLXL24, SLX20, WZC^{+}21]. heavy [SML22, ZMZ22]. heavy-tailed [SML22, ZMZ22]. Herriot [AS22, IK21]. heterogeneous [yCjL23, CIO24, LZ20b, LL23]. heteroscedastic [ASN22, Ros20, ZZC24]. heteroskedasticity [ST21]. hidden [Sab21, WT24]. hierarchical [Wal23, WT24]. High [FS23, HCX23, Kim20, YH23, BM21, CSP21, CG22, DJPZ22, FHL24, GC21, HS24a, JHH21, KU22, LX23, LZ22, MWL24, WY22, WL23, WZSQ24, WC24, WL20, XLZ20, XZ21, ZHB20, ZZGZ21]. High-dimensional [HCX23, Kim20, YH23, CSP21, CG22, DJPZ22, GC21, HS24a, JHH21, LX23, LZ22, WZSQ24, WC24, WL20, XLZ20, XZ21, ZHB20, ZZGZ21]. high-separation [HS24a]. Higher [PFM21, Gór24, Su22]. Higher-order [PFM21]. Hilbert [DR24, LLZL21]. history [EGC20]. Hochberg [CS20]. homogeneity [HXZZ20, LFLL21]. Horvitz [DS20b]. household [WZL^{+}22]. Huber [WZSQ24]. Hurst [SST20]. hypercube [WCL22, ZYL20]. hypotheses [NSC21, NS24, NBD21, PX24, SBP22]. Hypothesis [CF24, Khm21, NRP20]. Hyvärinen [CMMD21].

Identifiability [CB20]. identified [BCKL20]. identify [AFMY20]. identity [LX23]. ignoring [YW20]. II [KN24, YYYW21]. image [Wal23]. imperfect [SMB20]. implementation [MLM21]. important [SZ23, ZZ20]. Improved [XMO21]. improvement [Hel23]. improving [YM24]. imputation [GC20, ZZ24]. INAR [DT24]. INARp [BS21b]. incidence [AKE22]. inclusion [LAYT22]. incomplete [KVH^{+}23]. Incorporating [LAYT22]. Increasing [LW22, EMM21, LAO22]. independence [CBL20, TL24, ZLHZ22]. independent [YZW23]. index [FTHL21, PPTO20, WL23, WC24, YYW23, ZFZ20]. indicators [ZFZ20]. indices [Che24]. induced [KZ21]. inequalities [BCKL20]. inequality [FE23, LFLL21, XMO21, ZZWZ22]. inequality-constrained [XMO21]. inessential [PS21]. Inference [BCKL20, CBB24, GPG20a, WY24, AHB21, AGB20, AGNUC23, BM22, FMT21, GR24, HLLW24, HDLF20, HLDB22, KCH21, KU24, KCD23, LZ20b, LYH22, RKvS24, Rez22, RSV20, Sae21, Sae22, WM20, WWZZ20, WK23, WC24, XDZL20]. inferential [LJ20]. infinite [Gór24]. inflated [GC20, HH24, LPL21, SZ21]. influence [DCDS21]. Information [CWY20, TL24, FLL^{+}24, HK22, HB21, Lan23, Ove22, TPZ22]. Information-based [CWY20]. informative [LFR20]. INGARCH [SZ22]. innovation [BS21b]. inputs [MM20]. insurance [KKSA22]. integer [DK21, SWS24, YWYL20, ZLW^{+}24]. integer-valued [DK21, SWS24, YWYL20, ZLW^{+}24]. integrated [DR24, YX23]. integrative [Sab21]. intensities [BAA24]. interference [KZZ23]. interpolation [YCL21]. interval [LJ20, SLX20, TN23]. interval-censored [SLX20]. intervals [DJPZ22, HS24b, Jeo23, KWW20, KC24, LP21, PFM21]. intervention [HC23]. interventions [ZCQC21]. Invariant [AG21, DZ24]. inverse [BFMS21, MP23, SD21]. involves [LW20]. IPW [ZZ21]. IPW-based [ZZ21]. irregularly [EFR20]. isomorphic [LTZF22]. isotropy [WH20]. issues [GS24]. item [FHS23]. iterated [Kal23]. Iterative [AGB20].

J [GPG20a]. Jackknife [HS24b, SZ21, NMP21]. January [Ano20h, Ano21h, Ano22h, Ano23j, Ano24i]. Joint [LZ20b, MSFT23]. July [Ano20l, Ano21j, Ano22k, Ano23k, Ano24f]. jump [AG21, Buo20, Mie21, TN23]. jump-diffusion [AG21]. jumps [Buo20, MU21].

Kaplan [BI22]. Kernel [CMZ20, Pen24, HCX23, Kak20, LCP22, LLZL21, LZLL24, EFR20, ZHLT20, ZZ24]. kernel-based [HCX23]. Kinematic [TS22]. knockoffs [BDL^{+}23, PX24]. Koopman [CLPX23].

L2 [YH23]. lack [PSS20]. lack-of-fit [PSS20]. Lagrangian [ZDW^{+}21]. Laguerre [Dus21]. LAMN [TN23]. large [Fro22, LPK^{+}21]. Lasso [BCKL20, DT24, DCS20, SW22]. late [KS22]. latent [LZ22]. Latin [BS21a, UB22, WCL22, ZYL20]. laws [KOV22]. learning [KW24, LFW22, LZLL24]. Least [BdFL20, CWS20]. left [Che24]. left-truncated [Che24]. level [FHS20, HLYZ20, HWS24, LMMJ21, Sab21, TH22, WM24, ZXS23]. Lévy [Buo20]. life [EGC20]. lifetimes [KN24]. likelihood [AS22, BL20, BB24, FMT21, FHL24, GK21, GL20, HS24b, HLDB22, Jeo23, LCKB24, LPL21, LZ20b, MG23, SZ21, SLX20, WL23, YFS22, ZLW^{+}24]. likelihood-based [Jeo23, LCKB24, ZLW^{+}24]. likelihood-ratio [YFS22]. Limit [KOV22, LYZ21, Oui21, RW23]. limits [JHH21, Llo22]. line [WZL^{+}22]. linear [CWS20, Chu23, CMMD21, Don21, FE23, GAC23, GMFY20, GK22, GPP23, GHRS20, HXZZ20, HCZG24, KU21, KZ21, KCH21, Lan23, LAYT22, LYZ21, LLZL21, LPK^{+}21, LFR20, PdOA22, SDSZ21, SW22, SLW24, TJ20, WY24, WC24, XDZL20, XS21, XLZL22, Yin20, YW20, ZYLY22, ZSS23, ZLHZ22, ZFZ20]. linear-quadratic [ZYLY22]. link [FC20]. linkage [GR24, SSdW22]. links [PPTO20]. Local [BI22, Chu23, BKP20, MP21, Nar21, Oui21, Wal23]. Locally [WLFL24, GZ23, HY20, HDLF20]. locally-stationary [GZ23]. location [KC24]. location-scale [KC24]. locations [BG22a]. log [BB24, NMP21]. log-concave [BB24]. log-periodogram [NMP21]. Logical [NLL20]. logistic [CWY20, FY22, HY20, WMZ20, WG20]. logit [Jeo22, RD21]. long [dSCBP23, CLT23, DR24]. longitudinal [FHL24, LY21, LYH22, SCCF24]. loss [MS21b, WZSQ24]. losses [BFMS21, MB24]. low [CKSW21, FHM22]. low-rank [CKSW21]. lower [KLCZ23, KH20, SZ23]. lower-order [KLCZ23, SZ23].

m [KLCZ23]. machine [ZDW^{+}21]. main [FHS20, Gro20]. maintenance [SMB20]. Malliavin [SST20]. Mallows [ZZ24]. manifold [KLR23]. Mann [AHZ20]. MANOVA [SBP22, YH23]. many [BCKL20]. March [Ano20k, Ano21l, Ano22i, Ano23l, Ano24g]. Marginal [LB22, NBD21]. marginals [MSFT23]. Markov [CF24, Cav22, CBL20, FC20, Khm21, LZ20b, Pen24, WT24, ZDW^{+}21]. Markovian [BAA24, Mie21]. Martingale [XC20]. Martingale-difference-divergence-based [XC20]. mass [TTML21]. massive [QLXL24, YYW23]. matching [JT23, VdC21]. matrices [BFMS21, HWS24, LYZ21, RKvS24, SBP22, WL20, YK23, ZHB20]. Matrix [DGCL21, BM21, BFMS21, CKSW21, CSP21, FY22, JHH21, KZ21, LX23, LZW^{+}24, WZSQ24, YX23, YCL21, ZYG22]. matrix-variate [FY22]. max [Wal23]. max-pooling [Wal23]. maximin [ZYL20]. maximizing [MSFT23]. Maximum [BB24, GK21, XZ21, YJL24, AS22, BL20, FMT21, SLX20, Zha21]. Maximum-type [XZ21]. MaxT [NKST21]. May [Ano20j, Ano21i, Ano22g, Ano23h, Ano24h]. MCMC [MAC22]. Mean [WL24, dSCBP23, BKM21, CSP21, FHS23, HK22, HCX23, LP21, MRS21, SZ21, TJ20, XMO21, YZZ22, YK23]. mean-covariance [YZZ22]. meaning [Gro20]. means [TJ20]. measure [HI22, PdOA22, ST21, ZY22]. measurement [HCZG24]. Measures [FZ24, DCDS21]. mechanically [BNF^{+}23]. Meier [BI22]. membership [ZHLT20]. memory [CLT23, DR24]. merged [Sae22]. method [AS22, BB21, Cla21, Jeo23, KU22, LPL21, LPK^{+}21, LYCL24, XMX21, Yan20, YCL21]. methodology [EGC20]. methods [DLZ23, HC23, LCKB24, QLZ23, XLZ20]. metric [Sal21]. mid [CS20]. Minimax [GZ20, LZLL24, MRS21, Kim20]. minimaxity [MS21a]. minimizing [OYF20]. minimum [GK22, KLCZ23, SZ23, ZZ20]. misclassification [FY22]. misclassified [CBB24]. missing [GK21, LPL21, LYH20, LY21, WL23, YZZ22, ZHLT20, ZZ24, ZFZ20]. misspecification [XS21]. misspecified [AG20, DJPZ22]. Mixed [HLYZ20, DJPZ22, GPP23, HWS24, KZ21, PdOA22, Pru22, SLJ20, SCCF24, TJ20, WY24, XS21, YW20]. mixed-level [HWS24]. mixture [BB24, BFMS21, CCHR22, LZ20a, MB20, SWS24, TTML21, WZC^{+}21]. mixtures [MS23]. MLE [YM24]. Model [DZ24, WM20, YYW23, YZW23, Yin20, AG20, ASN22, AS22, BJZ22, dSCBP23, BS21b, BL22, CSL20, CCHR22, CBB24, CQY22, CLT23, DJPZ22, FTHL21, FE23, GMFY20, HK22, Has21, HH21, HXZZ20, HCZG24, IK21, Jaz24, KWW20, KU21, KZ21, KN24, KCH21, LPL21, LB22, LTW21, LYH22, OYF20, PPTO20, QLXL24, QW23, Rez22, SDSZ21, SLJ20, SLW24, SLX20, TJ20, TLT22, Wal23, WZC^{+}21, WZL^{+}22, WLFL24, WF20, XS21, Yan20, ZWYX20, ZC20a, ZYG22, ZYLY22, ZSS23, ZY22, ZZWZ22, ZZ24]. Model-free [DZ24, WM20, Yin20]. modeling [BNF^{+}23, LNH^{+}23]. Modelling [Kal23, KKH24]. Models [LCZ24, AGB20, BCKL20, Cav22, CB20, CWS20, CMMD21, CBL20, DMK22, DGCL21, DGCL22, EGC20, FDGD24, FMT21, FHS20, GZ20, GAC23, GK22, GPP23, GS24, GC21, GHRS20, GL20, Hai22, HLLW24, HY20, HH24, HYD24, HDLF20, HCZG24, HHL20, Jeo22, JK20, KvW23, KZZ23, Lan23, LAYT22, LCP22, LYH20, LPK^{+}21, LFR20, LY21, LZ22, LL23, LJ20, LY20, LW22, MG23, PSS20, PdOA22, Pru22, Ros20, SML22, SA20, SW22, WMZ20, WY24, WT24, WC24, XC20, XLZ20, XS21, XLZL22, YYW23, YJL24, YZZ22, YW20, YFS22, ZLW^{+}24, ZLHZ22, ZFZ20]. Moderate [Pro20]. moderately [AP21]. modification [NKST21]. Modified [ZHB20, BAA24, Hel23]. Moment [CLT23, BCKL20, LYH20, Pil20]. Monitoring [JK20, GZZ20]. monotone [BdFL20, Buo20, DZ24, PdOA22]. Monte [MM20]. Morris [BFMS21]. most [Rya22]. Moving [TW20, DR24]. multi [AHY21, GZ20, LWL^{+}24]. multi-armed [AHY21]. multi-dimensional [LWL^{+}24]. multi-factors [GZ20]. multidimensional [AVVW24]. multilevel [KN24]. Multimatricvariate [DGCL22]. Multinomial [Kim20, HY20, Oui21, ZWYX20]. Multiple [WT24, ACCY21, BB21, CLPX23, GHRS20, HHL20, LZLL24, NLL20, PPTO20, Pru22, Sae22, WD23, XDZL20, ZCQC21]. multiple-design [Pru22]. multiple-links [PPTO20]. multistage [NRP20]. multivalued [AHZ20]. Multivariate [AGNUC23, Kak20, AS22, BG22a, BM22, DSP23, Dus21, GZ20, GKO22, IK21, KCH21, LCYZ21, MP21, Nar21, SSC20, YX23, Yin20].

Nadaraya [EFR20]. nature [DHS21]. Nearly [BP21, LLY21, Pro20]. negative [HH24, SWS24]. negligible [DHS21]. nested [BFR21, LW22]. net [Hel23, LAYT22, Pil20]. network [Wal23]. networks [GK21, KLR23, LWL^{+}24]. neural [KLR23, LWL^{+}24, Wal23]. noise [BBS21, CLT23, BBCS24, Pil20, ZMZ22]. noises [SML22]. noisy [LWL^{+}24]. non [AHY21, AP21, DHS21, Has21, Kak20, LFR20, LTZF22, MB24, WZC^{+}21]. non-central [Kak20]. non-curable [WZC^{+}21]. non-isomorphic [LTZF22]. non-linear [LFR20]. non-negligible [DHS21]. non-parametric [AHY21, MB24]. non-regular [Has21]. non-stationary [AP21]. nonignorable [LY21]. nonlinear [FMT21, LYH22]. nonlinearity [FZ24]. nonnegative [FDGD24, HHL20, Kak20]. Nonparametric [CF20, HC23, MT23, Sae21, AFMY20, BG22a, BKM21, CCHR22, ST21, WST21, WH20, BS21b]. nonparametrically [Chu23]. nonparametrics [BG22b]. nonparanormal [MG23]. nonregular [CT22b]. nonstationary [KCD23]. norm [DT22, FE23, LX23, YH23]. normal [BB24, FMT21, HK22, HXZZ20, JX22, MB20, TTML21, YK23, ZZGZ21, WL24]. normality [KRX21, MSR21]. Normalized [YHDB22, MU21]. norming [LFR20]. norms [KOV22]. Note [Ano24j, DCS20, FHM22, KVH^{+}23, Ste22]. novel [YM24]. nuisance [Has21, Hil24]. null [NBD21, PX24, YH23]. number [HB21, Hua21, KH20, Rya22]. numerical [Jeo23].

objective [AVVW24]. objects [ASN22]. observation [HK22]. observational [GR24]. observations [GK21, KU21, KU24, TN23, WL23, YZZ22]. observed [Kal23, Lan23]. obtain [Jeo23]. odds [LAO22, Tan22]. ODE [LWL^{+}24]. one [MRS21, NSC21, ZZC24]. one- [NSC21]. one-parameter [MRS21]. one-way [ZZC24]. online [PW21]. operating [BFVT22, WM20, WK23]. operator [DT22]. operators [DR24]. Optimal [AHB21, BJZ22, CKSW21, DMK22, FHS20, FHS23, GAC23, HYD24, HCZG24, HHL20, KKSA22, LLZL21, LMMJ21, NMP21, NRP20, QLXL24, Ros20, SS20, WVJ^{+}20, YZZ22, ZYL20, CWY20, Cla21, GZ20, GPP23, Hai22, HY20, HB21, HH21, JT23, LZ20a, LYCL24, PS21, SKS21, Tsa22, WY22, YYYW21, YZW23, ZC20b]. Optimality [GHRS20, Kim20]. optimization [Don21]. optimizing [OYF20]. optional [TLG24]. oracle [FE23]. Orbit [FC20]. Orbit-based [FC20]. ORDANOVA [GM21]. Order [SWS24, GPG20a, GPG20b, KLCZ23, PFM21, SZ23, Tsa22, WM22, YWYL20, Zha21]. order-of-addition [Tsa22, WM22]. ordering [Sal21]. ordinal [GM21]. ordinary [Cla21]. Ornstein [Kal23]. Orthogonal [ZLZ20, BCF^{+}24, CT22a, Hua21, LLY21, LCYZ21, LTZF22, LMMJ21, WYL21, WY22]. orthogonally [BM21]. Outcome [ZY22, LNH^{+}23]. Outcome-adjusted [ZY22]. outcomes [LC20]. overall [PL20]. Overlap [yCjL23]. overlapping [SM23].

p [DT24]. Pages [Ano20i, Ano20l, Ano20k, Ano20j, Ano20g, Ano21k, Ano21h, Ano21j, Ano21l, Ano21i, Ano21g, Ano22j, Ano22h, Ano22k, Ano22i, Ano22g, Ano22l, Ano23i, Ano23j, Ano23k, Ano23l, Ano23h, Ano23g, Ano24i, Ano20h]. pair [YM24]. paired [GD20, Gro20]. pairwise [BS21a, NLL20]. pairwise-variance [BS21a]. panel [WWZZ20]. Panels [CF24]. parabolic [KU21]. parallel [LTW21, WM24]. parameter [Cla21, GKO22, Has21, Hil24, Jeo23, JK20, Kal23, MRS21, NMP21, SML22, SST20]. parameterization [LLT22]. parameters [DHS21, GZ23, HH21, LB22, OYF20, YJL24]. Parametric [KU21, AHY21, BB24, CF24, LFLL21, MB24, YYW23]. partial [HXZZ20, WK23, ZSS23]. partially [BCKL20, FE23, GMFY20, HCZG24, KCH21, WC24, XDZL20, ZFZ20]. Penalised [MAC22, MMD20]. Penalized [CM23, LCP22, FHL24, KV21]. penalty [WLFL24]. percentile [LNH^{+}23]. percentiles [KC24]. performance [ZDW^{+}21]. periodic [BS21b, MT23]. periodogram [MMD20, NMP21]. permutation [ZXS23]. permutations [FC20]. perspectives [BDL^{+}23]. perturbed [Baz23]. pFDR [HB21]. phase [CIO24, YYYW21]. Piecewise [DK21]. Pillai [ZHB20]. Plann [GPG20a]. plans [Gro20]. planted [KvW23]. plausibility [BL22]. plot [HLYZ20, SZ23, ZZ20]. plugin [GZ23]. pmf [BdFL20]. point [HY20, LG21, LCYZ21, LZ22, TTML21, ZH23, ZYLY22]. point-mass [TTML21]. points [KS22, KZA23, PS21]. Poisson [MU21, RW23, SS20]. polynomial [BI22, Hai22, MP21]. polynomials [LZ20a]. pooling [Wal23]. populations [CIO24, SZ21]. Portmanteau [ZMZ22]. Portmanteau-type [ZMZ22]. positive [RKvS24]. possible [XS21]. possibly [KY24]. post [GAC23, HC23]. post-selection [GAC23]. Posterior [AG20, Jeo22, BM22, LB22, WM20]. power [CQY22, YHDB22]. powerful [Rya22, WL24]. pre [HC23]. pre-post [HC23]. precise [Oui21]. precision [LZW^{+}24, YX23, ZYG22]. prediction [LLZL21, LJ20, TJ20]. Predictive [MS21b, HK22]. predictor [SLJ20, TJ20]. presence [BS21c, Has21, MT23]. Preserving [GC20, TH22]. price [TPZ22]. Principal [ZC20a, AP21, RX20]. prior [AVVW24, HB21, MSFT23, YHDB22]. priors [BS22, Has21, LB22, MS23, RD21, WG20, YM24, ZYG22]. probabilities [Baz23, LAYT22, YZW23]. probability [MSFT23, Tan22]. problem [AHY21, ASN22, Buo20, Cla21, LFLL21, VdC21]. problems [Pru22, YX23, ZZGZ21]. procedure [ACCY21, BB21, CS20, NKST21, Rez22]. procedures [ACCY21, DT22]. process [AGB20, AM23, Baz23, BBCS24, KCD23, LB22, SWS24, SSC20]. Processes [CF24, AG21, BL20, BP21, BOAB20, Buo20, DT24, DR24, FHM22, Kal23, KU22, KW24, KU24, Nar21, Pro20, SSC20, SZ22, TN23, YWYL20, ZWL22, ZC20b]. product [RX20]. Projected [WL20]. Projection [KRX21, Dus21, LZLL24, TH22, ZXS23]. projections [Sab21]. propensity [yCjL23, WXT^{+}21, ZY22]. Properties [AGNUC23, LAO22, Ove22, BL20, BOAB20, HI22, KWW20, LCZ24, TH22, WG20]. property [TN23]. proportion [NBD21]. proportional [HH21, QLXL24, WZC^{+}21]. provides [KKSA22]. proving [MS21a]. Proximal [DT22]. Pseudo [WY20]. Publishers [Ano24j]. purely [Buo20]. pursuit [KRX21].

quadratic [LZ20a, WWZZ20, ZYLY22]. quantification [KvW23]. Quantile [WWZZ20, ZFZ20, LCP22, LLZL21, MMD20, SDSZ21, TW20, WL23, WLFL24, XC20]. quantiles [BM22]. quantum [CKSW21]. Quasi [GL20, ZH23, NZ22, ZYG22]. Quasi-Bayesian [ZH23, ZYG22]. Quasi-likelihood [GL20]. quasi-random [NZ22].

Random [LFR20, LZ20b, NZ22, Yan20, BFR21, GHRS20, KZ21, LYH20, LJ20, LZLL24, MB20, MM20, Rya22, YWYL20, ZHLT20, ZZ24]. randomized [GZZ20, JS22, Lop20, ZCQC21]. randomness [HI22]. range [dSCBP23, ZWYX20]. rank [CKSW21, MWL24, MG23]. rank-likelihood [MG23]. rate [EMM21, GZ23, HLDB22, LAO22, NSC21, NS24, ZC20b]. rates [LZLL24]. ratio [KC24, YFS22]. ratios [Tan22]. re [GZZ20]. re-estimation [GZZ20]. reach [CFM23]. Read [RW23]. receiver [WM20, WK23]. reclassification [Hel23]. record [GR24]. recovering [KZA23]. recovery [HLDB22]. rectangles [UB22]. recurrent [CSL20, SA20]. reduced [MWL24]. reduction [Don21, DSP23, EGC20, SD21]. Reference [Has21, ZZGZ21]. region [FHS20, IK21]. regions [GHRS20]. Regression [CB20, FDGD24, GR24, SSC20, SCCF24, AG20, CBB24, CMZ20, CWS20, CWY20, Chu23, DCS20, DSP23, FY22, FE23, GZ20, GK22, GHRS20, HY20, HH24, HYD24, KKH24, KLR23, LPL21, LCP22, LLZL21, LPK^{+}21, LB22, LFR20, LTW21, LNH^{+}23, LW22, MWL24, EFR20, MP21, OYF20, PSS20, RD21, SDSZ21, SD21, TW20, WZL20, WWZZ20, WZL^{+}22, WL23, WLFL24, WG20, WY20, XDZL20, XZ21, YX23, YJL24, YZW23, Yin20, ZZ24, ZFZ20].

Regression-assisted [GR24]. regressions [BJZ22]. regular [Has21, KLCZ23, TH22, UB22]. regular-graph [UB22]. Regularization [XLZ20, Cla21]. Regularized [FY22]. related [MSFT23]. relationship [Gro20]. Removing [PS21]. renewal [Baz23]. repeated [FHS23]. replacement [BKM21, LYCL24]. Representative [KZA23]. reproducing [LLZL21]. Rerandomization [JS22, BS21c, KKSA22]. Resampling [KY24]. residual [yCjL23]. residuals [MU21]. resolution [KH20]. response [CBB24, DSP23, FY22, FHS23, GZZ20, HHL20, LC20, MCJ23, Yan20, YYYW21]. response-adaptive [GZZ20, LC20, MCJ23]. responses [Jeo22, LFW22, MCJ23, XLZL22, ZZ24]. Restricted [AKE22, FHS20, MRS21]. result [Llo22]. Results [KLCZ23, DSK21, MS23, MSFT23, NS24]. Review [MS21a, yCjL23, CBL20, Don21, DSP23, MS23]. Ridge [BS21c, OYF20]. right [Che24]. right-censored [Che24]. risk [Baz23, Jaz24, WZC^{+}21]. risks [AKE22]. Robust [FHL24, HLLW24, LYH20, LFW22, RSV20, WST21, XS21, LP21, LGM23, MWL24, WZSQ24, ZHLT20]. root [ZMZ22]. row [KVH^{+}23]. row-column [KVH^{+}23]. ruin [Baz23]. rule [AVVW24, CMMD21]. run [Tsa22].

sample [BM21, BKM21, GZZ20, GAC23, HLP23, JHH21, KWW20, LYZ21, LFLL21, MPSV22, SM23, Tan22, TLG24, XLZL22, ZHB20, ZZGZ21]. sampled [FHM22]. samples [Gór24, KOV22, Sae21]. sampling [DS20b, JT23, Jaz24, LCKB24, MM20, SZ22, ZDW^{+}21]. Sanov [MP23]. Saturated [WM22, CT22a, DHS21]. Saunders [DGCL21, Kak20]. scalable [LPK^{+}21]. scalar [GS24, ZSS23]. Scale [KN24, BFMS21, KV21, KC24, LPK^{+}21, MS23, YJL24]. scan [ACCY21]. scatter [BFMS21]. score [yCjL23, WXT^{+}21, ZY22]. score-based [WXT^{+}21]. scores [XZ21]. scoring [AVVW24, CMMD21]. screening [Che24, DZ24, HCX23, HWS24, PL20, WMZ20]. searching [GPP23]. seasonal [BNF^{+}23, MR21]. second [Pil20, Zha21]. second-order [Zha21]. section [KvW23]. seemingly [WY24]. selection [CSL20, CQY22, CWY20, CM20, GAC23, GC21, LAYT22, LFR20, LTW21, LL23, OYF20, PX24, SW22, WY20, ZY22, ZFZ20]. selective [DSP23]. self [MU21]. self-normalized [MU21]. Semi [CF24, LFLL21, BS21a, BB24, UB22]. semi-Latin [BS21a, UB22]. Semi-Markov [CF24]. Semi-parametric [LFLL21, BB24]. semimartingales [Mie21]. Semiparametric [LNH^{+}23, LGM23, MPSV22, Sae22, SMB20, SLX20, WZC^{+}21, WXT^{+}21, HANB20, HCZG24, SA20]. sensitive [LTW21]. separable [LYZ21]. separation [HS24a]. September [Ano20g, Ano21g, Ano22l, Ano23g, Ano24k]. sequence [Khm21, MS21a, MT23, WST21, WF20]. Sequential [GZZ20, PW21, CLPX23, GC21, HB21, KS22, LFR20, LC20, Llo22, NRP20, TLG24, ZCQC21]. series [AP21, BAA24, dSCBP23, BM20, DT24, CMMD21, DK21, GZ23, LCYZ21, LCZ24, LY21, LZ22, LY20, Yin20, ZWYX20, Zha21]. set [Jaz24]. sets [ZZ20]. setting [CG22]. settings [CSP21]. sharp [Lop20]. shortfall [MB24]. shrinkage [BS22, WG20, ZYG22]. sieve [SLX20]. sign [MLM21]. signal [BS22]. signals [HB21]. SiMaFlex [BB21]. simple [CWS20]. simplex [HH21]. simulation [FS23]. simulations [MM20]. Simultaneous [KCH21, TS22, WK23, CSP21]. single [FTHL21, PPTO20, WL23, WC24, YYW23, ZFZ20]. single-index [FTHL21, PPTO20, WL23, WC24, YYW23, ZFZ20]. size [GZZ20, LCKB24, LFLL21, LW22, UB22]. size-biased [LCKB24]. sizes [BM21, Tsa22]. Skew [HXZZ20, JX22]. Skew-normal [HXZZ20, JX22]. skewed [SZ21]. skewness [JX22]. slab [LAYT22]. Sliced [WCL22]. SLOPE [DT22]. Small [PL20, BM21, Gór24, TJ20]. Smooth [LY20, KY24, MT23]. smoothed [Hil24]. smoothing [BI22, GKO22, WVJ^{+}20]. smoothness [Ste22]. sn [KLCZ23]. sn-m [KLCZ23]. softmax [YZW23]. sojourn [CF24]. solution [AHY21]. some [DMK22, NS24, Oui21]. sorted [DT22]. sources [Sae22]. space [CT22b, DR24, Rez22, WYL21]. space-filling [CT22b, WYL21]. space-valued [DR24]. spaced [EFR20]. spaces [LLZL21, Sal21]. spacings [SM23]. Sparse [BFR21, LZLL24, CKSW21, DCS20, GK21, Jeo22, KvW23, LPK^{+}21, WLFL24, XLZ20]. spatial [LAYT22, EFR20, Su22, TJ20, WH20]. spatiotemporal [HH24, WD23]. SPDE [KU21]. specific [BB21]. spectral [Cav22, LLBR22, LYZ21, MMD20, WH20]. spherical [Su22]. spike [LAYT22]. spike-and-slab [LAYT22]. spiked [JHH21]. spline [CM23, LZ20a, WXT^{+}21, Yan20]. splines [KV21, RX20, WVJ^{+}20]. split [HLYZ20, SZ23, ZZ20]. split-plot [HLYZ20, SZ23, ZZ20]. spread [GR24]. spurious [NKST21]. squared [HK22, TJ20]. squares [BS21a, BdFL20, CWS20]. stable [CLPX23, FHM22, Pro20]. Stacy [AM23]. standard [BB24]. standardization [DCS20]. state [Mie21, Rez22]. state-dependent [Mie21]. state-space [Rez22]. stationarity [GS24, Zha21]. stationary [AP21, GZ23, HDLF20]. Statist [GPG20a]. statistic [FHBD20, Hel23, SM23, YH23, ZLHZ22]. Statistical [ASN22, HDLF20, KU24, RKvS24, JT23, LAO22, LYH22]. Statistics [BdFL20, BG22b, HLP23, Kim20, KS22, LYZ21, RW23, ZHB20, Zha21]. step [KN24]. step-stress [KN24]. stepped [TL24]. stick [RD21]. stick-breaking [RD21]. stochastic [BP21, Gór24, MB20, XMX21, YFS22, ZC20b]. stopping [TLG24]. strategy [BS21c]. stratification [JS22]. Stratified [MM20, Sae21]. strength [WY22]. stress [KN24]. Strong [CWS20, BNF^{+}23, LLY21]. structure [BCF^{+}24, dSCBP23, GL20, LAYT22, NS24]. structure-changed [GL20]. structures [NSC21]. studies [DJPZ22, GR24, QW23, WT24]. subdata [CWY20]. Subgroup [SLW24, HLLW24]. subject [FY22]. subsampling [PW21, QLXL24, YZW23]. subset [WL24, YYYW21]. substitute [JS22]. sufficient [Don21, DSP23, SD21]. sums [FDGD24, Fro22]. sup [FE23]. sup-norm [FE23]. supersaturated [LYCL24, WM22]. Support [HY20, ZDW^{+}21]. surrogate [AGB20, LPK^{+}21]. survey [JT23, XMO21]. surveys [GC20, LY21, LTW21]. survival [Che24, LC20, MCJ23, QLXL24, Tan22]. Switching [Cav22]. SYMARFIMA [dSCBP23]. symmetric [dSCBP23, MPSV22, RKvS24]. symmetrical [WCL22]. synchronous [SCCF24]. system [SMB20]. systematic [LLT22, SZ22]. systems [CKSW21].

tables [SSdW22, Tan22]. tail [GKO22]. tailed [SML22, ZMZ22]. tale [SSdW22]. tapered [Nar21]. Targeted [GPG20a, GPG20b]. techniques [KY24]. Temporal [SZ22]. tending [YJL24]. tending-to-zero [YJL24]. Tensor [RX20]. test [AHZ20, BBS21, BKM21, CSP21, CLPX23, Hil24, HXZZ20, LX23, LFLL21, MPSV22, NLL20, WH20, YFS22, Zha21, ZSS23, ZLW^{+}24, ZMZ22, ZZC24]. Testing [BM20, Gór24, JX22, Khm21, MR21, PSS20, SBP22, ACCY21, BB21, CF24, CB20, DT22, FS23, FHS23, GPG20a, GPG20b, KU22, NZ22, NRP20, PdOA22, WD23, WT24, YH23, ZLHZ22]. tests [BG22a, Cav22, FC20, KN24, KS22, KRX21, MSR21, MK21, Rya22, TLT22, TLG24, WL24, WL20, XC20, XZ21, Yin20]. Texas [GPG20a, GPG20b]. their [BS21a]. theorem [LYZ21, MP23, Oui21]. theorems [Pru22, RW23]. theory [Cla21, KCD23]. there [Hil24]. thinning [PW21, SWS24]. Thompson [DS20b]. three [Sab21]. three-level [Sab21]. threshold [ACCY21, DZ24, GL20, LW20, SWS24, ZYLY22]. threshold-type [ACCY21]. tight [WY22]. Time [GS24, AP21, AFMY20, BAA24, dSCBP23, BP21, BKM21, BM20, DT24, CF24, CMMD21, DK21, DSK21, FHS23, GZ23, LCYZ21, LCZ24, LW20, LY21, LZ22, LY20, QW23, TN23, ZWYX20, Zha21]. time-evolving [DSK21]. time-series [GZ23]. time-to-event [LW20]. times [KS22]. Tobit [WZL^{+}22]. total [AHB21, KZZ23, YYYW21]. trace [ZHB20]. Tractable [RD21]. Trading [TPZ22]. transformation [CQY22, HLLW24]. transition [Khm21, Pen24]. transport [JT23]. treatment [AHZ20, Ros20, ZZ21]. trend [JK20, MT23]. trial [BFVT22, HI22, Llo22, ZCQC21]. trials [GZZ20, YYYW21]. trigonometric [FDGD24]. true [NBD21]. truncated [Che24]. TV [WLFL24]. Two [BKM21, GM21, XLZL22, ZZGZ21, AKE22, CCHR22, FHS20, GR24, HLP23, KP24, LFLL21, LMMJ21, MPSV22, NSC21, SZ21, Tan22, TH22, UB22, WM24, ZHB20]. two-class [CCHR22]. two-level [FHS20, LMMJ21, TH22, WM24]. Two-sample [BKM21, XLZL22, ZZGZ21, HLP23, LFLL21, MPSV22, Tan22]. Two-way [GM21, NSC21]. type [ACCY21, BFMS21, BBCS24, KV21, XZ21, YFS22, ZMZ22, KN24]. Type-II [KN24]. types [KP24].

U [ZLHZ22]. U-statistic [ZLHZ22]. Uhlenbeck [Kal23]. ultrahigh [LZW^{+}24, LL23, WMZ20]. ultrahigh-dimensional [LL23, WMZ20]. unbiased [DHS21]. Uncertainty [BFVT22, KvW23]. Unified [LYH22, DS20a, ZLW^{+}24]. Uniform [BS21a, QLZ23, ZLZ20, ZXS23]. Uniformity [CT22a, MK21]. Uniformly [WL24, WC24]. unimodal [CM23]. unit [Hai22, ZMZ22]. univariate [CF20]. universal [Sal21]. Universally [CFM23]. unknown [HK22, KCH21, PL20, TTML21]. unrelated [WY24]. unstable [BP21, Pro20]. user [BB21]. Using [CT22b, BS22, BCKL20, Cav22, Dus21, HK22, HWS24, LB22, MG23, NKST21, SSC20, TS22, WXT^{+}21, XMO21, XZ21, ZWL22].

valid [WC24]. Validation [LPL21]. value [Hil24]. valued [BM20, DK21, DR24, LCZ24, SWS24, YWYL20, ZLW^{+}24]. values [CS20]. VAR [Cav22]. Variable [CQY22, PX24, CSL20, HCX23, LAYT22, LW20, LTW21, LJ20, WY20, ZFZ20]. variables [DCS20, GK22, HHL20, TLG24]. Variance [DJPZ22, BFR21, BS21a, DLZ23, EGC20, FHBD20, GPP23, HK22, PdOA22, PL20, XMO21, YCL21]. variance-covariance [YCL21]. variant [LTW21]. variate [DGCL21, FY22]. variation [GM21]. variations [SST20]. varying [FTHL21, LCP22]. vector [AP21, CSP21, SML22, ZDW^{+}21, WL24]. ventilated [BNF^{+}23]. via [Che24, Cla21, Has21, Jaz24, KKSA22, LWL^{+}24, LMMJ21, RD21, SST20, WST21, WT24, ZZ20, ZCQC21, ZXS23]. view [WK23]. volatility [Rez22, TPZ22].

Wagner [CM20]. walk [MAC22]. Wasserstein [MS21b]. Watson [EFR20]. wave [SST20]. wavelet [RKvS24]. way [GM21, NSC21, ZZC24]. weak [BBS21, CWS20]. weakly [Chu23, KW24]. wedge [TL24]. weight [yCjL23]. Weighted [GK22, HLP23]. weights [ZH23]. while [NSC21]. white [BBS21]. whiteness [BM20]. Whitney [AHZ20]. Whittle [BKP20, Nar21, SML22]. whole [SZ23, ZZ20]. whole-plot [SZ23]. wide [DJPZ22, WT24]. Wilcoxon [XZ21]. Wishart [BFMS21]. without [HB21]. Wolverton [CM20]. Wolverton-Wagner [CM20]. working [TL24].

Z [LCZ24]. Z-valued [LCZ24]. Zero [BAA24, GC20, HH24, LPL21, SZ21, YJL24]. zero-inflated [HH24, LPL21, SZ21]. Zero-modified [BAA24].

References

Arias-Castro:2021:SPM

[ACCY21] Ery Arias-Castro, Shiyun Chen, and Andrew Ying. A scan procedure for multiple testing: Beyond threshold-type procedures. Journal of Statistical Planning and Inference, 210(??):42–52, January 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

           (electronic). URL http://www.sciencedirect.com/science/
           article/pii/S037837582030046X.

Antonczyk:2020:NAI

[AFMY20] Dirk Antonczyk, Bernd Fitzenberger, Enno Mammen, and Kyusang Yu. A nonparametric approach to identify age, time, and cohort effects. Journal of Statistical Planning and Inference, 204(??): 96–115, January 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837581930045X.

Abraham:2020:PCM

[AG20] Christophe Abraham and Paul-Marie Grollemund. Posterior concentration for a misspecified Bayesian regression model with functional covariates. Journal of Statistical Planning and Inference, 208(??):58–65, September 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375820300069.

Amorino:2021:IDA

[AG21] Chiara Amorino and Arnaud Gloter. Invariant density adaptive estimation for ergodic jump-diffusion processes over anisotropic classes. Journal of Statistical Planning and Inference, 213(??): 106–129, July 2021. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837582030121X.

Alawieh:2020:ICG

[AGB20] Leen Alawieh, Jonathan Goodman, and John B. Bell. Iterative construction of Gaussian process surrogate models for Bayesian inference. Journal of Statistical Planning and Inference, 207(??): 55–72, July 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837581930103X.

Arbel:2023:MEB

[AGNUC23] Julyan Arbel, Stéphane Girard, Hien Duy Nguyen, and Antoine Usseglio-Carleve. Multivariate expectile-based distribution: Properties, Bayesian inference, and applications. Journal of Statistical Planning and Inference, 225(??):146–170, July 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822001082.

Aboukhamseen:2021:OCD

[AHB21] Suja Aboukhamseen, Shahariar Huda, and Mausumi Bose. Optimal crossover designs for inference on total effects. Journal of Statistical Planning and Inference, 213(??):253–261, July 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820301361.

Ai:2021:NPS

[AHY21] Mingyao Ai, Yimin Huang, and Jun Yu. A non-parametric solution to the multi-armed bandit problem with covariates. Journal of Statistical Planning and Inference, 211(??):402–413, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300902.

Ai:2020:MWT

[AHZ20] Chunrong Ai, Lukang Huang, and Zheng Zhang. A Mann–Whitney test of distributional effects in a multivalued treatment. Journal of Statistical Planning and Inference, 209(??):85–100, December 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300318.

Al-Kandari:2022:REC

[AKE22] Noriah Al-Kandari and Hammou El Barmi. Restricted estimation of the cumulative incidence functions of two competing risks. Journal of Statistical Planning and Inference, 217(??):122– 140, March 2022. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375821000720.

Arfe:2023:GBB

[AM23] Andrea Arfè and Pietro Muliere. A general Bayesian bootstrap for censored data based on the beta-Stacy process. Journal of Statistical Planning and Inference, 222(??):241–251, January 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000647.

Anonymous:2020:EBa

[Ano20a] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 204(??):ii, January 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375819300588.

Anonymous:2020:EBb

[Ano20b] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 205(??):ii, March 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375819300850.

Anonymous:2020:EBc

[Ano20c] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 206(??):ii, May 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375819301077.

Anonymous:2020:EBd

[Ano20d] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 207(??):ii, July 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375820300112.

Anonymous:2020:EBe

[Ano20e] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 208(??):ii, September 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820300288.

Anonymous:2020:EBf

[Ano20f] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 209(??):ii, December 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820300537.

Anonymous:2020:PS

[Ano20g] Anonymous. Pages 1–130 (September 2020). Journal of Statistical Planning and Inference, 208(??):1–130, September 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2020:PJa

[Ano20h] Anonymous. Pages 1–216 (January 2020). Journal of Statistical Planning and Inference, 204(??):1–216, January 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2020:PD

[Ano20i] Anonymous. Pages 1–292 (December 2020). Journal of Statistical Planning and Inference, 209(??):1–292, December 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2020:PMb

[Ano20j] Anonymous. Pages 1–312 (May 2020). Journal of Statistical Planning and Inference, 206(??):1–312, May 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2020:PMa

[Ano20k] Anonymous. Pages 1–330 (March 2020). Journal of Statistical Planning and Inference, 205(??):1–330, March 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2020:PJb

[Ano20l] Anonymous. Pages 1–344 (July 2020). Journal of Statistical Planning and Inference, 207(??):1–344, July 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2021:EBa

[Ano21a] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 210(??):ii, January 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375820300690.

Anonymous:2021:EBb

[Ano21b] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 211(??):ii, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375820300951.

Anonymous:2021:EBc

[Ano21c] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 212(??):ii, May 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375820301294.

Anonymous:2021:EBd

[Ano21d] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 213(??):ii, July 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375821000112.

Anonymous:2021:EBe

[Ano21e] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 214(??):ii, September 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000306.

Anonymous:2021:EBf

[Ano21f] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 215(??):ii, December 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000550.

Anonymous:2021:PS

[Ano21g] Anonymous. Pages 1–150 (September 2021). Journal of Statistical Planning and Inference, 214(??):1–150, September 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2021:PJa

[Ano21h] Anonymous. Pages 1–160 (January 2021). Journal of Statistical Planning and Inference, 210(??):1–160, January 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2021:PMb

[Ano21i] Anonymous. Pages 1–226 (May 2021). Journal of Statistical Planning and Inference, 212(??):1–226, May 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2021:PJb

[Ano21j] Anonymous. Pages 1–292 (July 2021). Journal of Statistical Planning and Inference, 213(??):1–292, July 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2021:PD

[Ano21k] Anonymous. Pages 1–388 (December 2021). Journal of Statistical Planning and Inference, 215(??):1–388, December 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2021:PMa

[Ano21l] Anonymous. Pages 1–438 (March 2021). Journal of Statistical Planning and Inference, 211(??):1–438, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2022:EBa

[Ano22a] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 216(??):ii, January 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375821000859.

Anonymous:2022:EBb

[Ano22b] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 217(??):ii, March 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375821001002.

Anonymous:2022:EBc

[Ano22c] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 218(??):ii, May 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375821001178.

Anonymous:2022:EBd

[Ano22d] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 219(??):ii, July 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375822000088.

Anonymous:2022:EBe

[Ano22e] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 220(??):ii, September 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375822000258.

Anonymous:2022:EBf

[Ano22f] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 221(??):ii, December 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375822000568.

Anonymous:2022:PMb

[Ano22g] Anonymous. Pages 1–164 (May 2022). Journal of Statistical Planning and Inference, 218(??):1–164, May 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2022:PJa

[Ano22h] Anonymous. Pages 1–206 (January 2022). Journal of Statistical Planning and Inference, 216(??):1–206, January 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2022:PMa

[Ano22i] Anonymous. Pages 1–278 (March 2022). Journal of Statistical Planning and Inference, 217(??):1–278, March 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2022:PD

[Ano22j] Anonymous. Pages 1–326 (December 2022). Journal of Statistical Planning and Inference, 221(??):1–326, December 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2022:PJb

[Ano22k] Anonymous. Pages 1–332 (July 2022). Journal of Statistical Planning and Inference, 219(??):1–332, July 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2022:PS

[Ano22l] Anonymous. Pages 1–82 (September 2022). Journal of Statistical Planning and Inference, 220(??):1–82, September 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2023:EBa

[Ano23a] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 222(??):ii, January 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375822000738.

Anonymous:2023:EBb

[Ano23b] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 223(??):ii, March 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375822000866.

Anonymous:2023:EBc

[Ano23c] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 224(??):ii, May 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375822001057.

Anonymous:2023:EBd

[Ano23d] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 225(??):ii, July 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375823000101.

Anonymous:2023:EBe

[Ano23e] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 226(??):ii, September 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837582300023X.

Anonymous:2023:EBf

[Ano23f] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 227(??):ii, December 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375823000393.

Anonymous:2023:PS

[Ano23g] Anonymous. Pages 1–100 (September 2023). Journal of Statistical Planning and Inference, 226(??):1–100, September 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2023:PMb

[Ano23h] Anonymous. Pages 1–126 (May 2023). Journal of Statistical Planning and Inference, 224(??):1–126, May 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2023:PD

[Ano23i] Anonymous. Pages 1–162 (December 2023). Journal of Statistical Planning and Inference, 227(??):1–162, December 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2023:PJa

[Ano23j] Anonymous. Pages 1–272 (January 2023). Journal of Statistical Planning and Inference, 222(??):1–272, January 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2023:PJb

[Ano23k] Anonymous. Pages 1–282 (July 2023). Journal of Statistical Planning and Inference, 225(??):1–282, July 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2023:PMa

[Ano23l] Anonymous. Pages 1–74 (March 2023). Journal of Statistical Planning and Inference, 223(??):1–74, March 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2024:EBa

[Ano24a] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 228(??):ii, January 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375823000605.

Anonymous:2024:EBb

[Ano24b] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 229(??):Article 106106, March 2024. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375823000757.

Anonymous:2024:EBc

[Ano24c] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 230(??):Article 106127, May 2024. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375823000964.

Anonymous:2024:EBd

[Ano24d] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 231(??):Article 106157, July 2024. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375824000144.

Anonymous:2024:EBe

[Ano24e] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 232(??):Article 106178, September 2024. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375824000351.

Anonymous:2024:Ja

[Ano24f] Anonymous. July 2024. Journal of Statistical Planning and Inference, 231(??):??, July 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2024:Ma

[Ano24g] Anonymous. March 2024. Journal of Statistical Planning and Inference, 229(??):??, March 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2024:Mb

[Ano24h] Anonymous. May 2024. Journal of Statistical Planning and Inference, 230(??):??, May 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2024:PJa

[Ano24i] Anonymous. Pages 1–116 (January 2024). Journal of Statistical Planning and Inference, 228(??):1–116, January 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:2024:PN

[Ano24j] Anonymous. Publishers note. Journal of Statistical Planning and Inference, 229(??):Article 106107, March 2024. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375823000769.

Anonymous:2024:S

[Ano24k] Anonymous. September 2024. Journal of Statistical Planning and Inference, 232(??):??, September 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Alshammri:2021:GPC

[AP21] Fayed Alshammri and Jiazhu Pan. Generalized principal component analysis for moderately non-stationary vector time series. Journal of Statistical Planning and Inference, 212(??):201– 225, May 2021. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375820301130.

Angkunsit:2022:AML

[AS22] Annop Angkunsit and Jiraphan Suntornchost. Adjusted maximum likelihood method for multivariate Fay-herriot model. Journal of Statistical Planning and Inference, 219(??):231–249, July 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821001348.

Al-Sharadqah:2022:SAC

[ASN22] Ali Al-Sharadqah and Phuc Nguyen. Statistical analysis of concentric objects estimation problem under the heteroscedastic Berman model. Journal of Statistical Planning and Inference, 218(??): 9–24, May 2022. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375821000938.

Antoniano-Villalobos:2024:MOP

[AVVW24] Isadora Antoniano-Villalobos, Cristiano Villa, and Stephen G. Walker. A multidimensional objective prior distribution from a scoring rule. Journal of Statistical Planning and Inference, 231(??):Article 106122, July 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375823000915.

Balakrishna:2024:ZMC

[BAA24] N. Balakrishna, P. Muhammed Anvar, and Bovas Abraham. Zeromodified count time series with Markovian intensities. Journal of Statistical Planning and Inference, 229(??):Article 106092, March 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823000538.

Bazyari:2023:RPG

[Baz23] Abouzar Bazyari. On the ruin probabilities for a general perturbed renewal risk process. Journal of Statistical Planning and Inference, 227(??):1–17, December 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375823000174.

Bartenschlager:2021:NUS

[BB21] Christina C. Bartenschlager and Jens O. Brunner. A new user specific multiple testing method for business applications: the SiMaFlex procedure. Journal of Statistical Planning and Inference, 214(??):25–40, September 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000057.

Balabdaoui:2024:MLE

[BB24] Fadoua Balabdaoui and Harald Besdziek. Maximum likelihood estimation of the log-concave component in a semi-parametric mixture with a standard normal density. Journal of Statistical Planning and Inference, 230(??):Article 106113, May 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823000824.

Hariz:2024:FAE

[BBCS24] Samir Ben Hariz, Alexandre Brouste, Chunhao Cai, and Marius Soltane. Fast and asymptotically-efficient estimation in an autoregressive process with fractional type noise. Journal of Statistical Planning and Inference, 232(??):Article 106148, September 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375824000053.

Bhattacharjee:2021:WNT

[BBS21] Monika Bhattacharjee, Arup Bose, and Radhendushka Srivastava. A white noise test under weak conditions. Journal of Statistical Planning and Inference, 211(??):362–390, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300860.

Bailey:2024:DHD

[BCF^{+}24] R. A. Bailey, Peter J. Cameron, Dário Ferreira, Sandra S. Ferreira, and Célia Nunes. Designs for half-diallel experiments with commutative orthogonal block structure. Journal of Statistical Planning and Inference, 231(??):Article 106139, July 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823001088.

Bugni:2020:IPI

[BCKL20] Federico A. Bugni, Mehmet Caner, Anders Bredahl Kock, and Soumendra Lahiri. Inference in partially identified models with many moment inequalities using lasso. Journal of Statistical Planning and Inference, 206(??):211–248, May 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375819300941.

Balabdaoui:2020:LSE

[BdFL20] Fadoua Balabdaoui and Gabriella de Fournas-Labrosse. Least squares estimation of a completely monotone pmf: From analysis to statistics. Journal of Statistical Planning and Inference, 204(??): 55–71, January 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375818302064.

Berti:2023:NPK

[BDL^{+}23] Patrizia Berti, Emanuela Dreassi, Fabrizio Leisen, Luca Pratelli, and Pietro Rigo. New perspectives on knockoffs construction. Journal of Statistical Planning and Inference, 223(??):1–14, March 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000696.

Boukehil:2021:EIS

[BFMS21] Djamila Boukehil, Dominique Fourdrinier, Fatiha Mezoued, and William E. Strawderman. Estimation of the inverse scatter matrix for a scale mixture of Wishart matrices under Efron–Morris type losses. Journal of Statistical Planning and Inference, 215(??):368– 387, December 2021. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375821000409.

Bailey:2021:SDE

[BFR21] R. A. Bailey, Célia Fernandes, and Paulo Ramos. Sparse designs for estimating variance components of nested factors with random effects. Journal of Statistical Planning and Inference, 214(??):76– 88, September 2021. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375821000033.

Bonsaglio:2022:AOC

[BFVT22] Marta Bonsaglio, Sandra Fortini, Steffen Ventz, and Lorenzo Trippa. Approximating the operating characteristics of Bayesian uncertainty directed trial designs. Journal of Statistical Planning and Inference, 221(??):90–99, December 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375822000283.

Bhattacharya:2022:BNT

[BG22a] Indrabati Bhattacharya and Subhashis Ghosal. Bayesian nonparametric tests for multivariate locations. Journal of Statistical Planning and Inference, 219(??):1–12, July 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821001075.

Binette:2022:BND

[BG22b] Olivier Binette and Simon Guillotte. Bayesian nonparametrics for directional statistics. Journal of Statistical Planning and Inference, 216(??):118–134, January 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000586.

Bagkavos:2022:LPS

[BI22] Dimitrios Bagkavos and Dimitrios Ioannides. Local polynomial smoothing based on the Kaplan–Meier estimate. Journal of Statistical Planning and Inference, 221(??):212–229, December 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000362.

Bai:2022:OMA

[BJZ22] Yang Bai, Rongjie Jiang, and Mengli Zhang. Optimal model averaging estimator for expectile regressions. Journal of Statistical Planning and Inference, 217(??):204–223, March 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000902.

Bhattacharyya:2021:TSN

[BKM21] Dhrubasish Bhattacharyya, Ruhul Ali Khan, and Murari Mitra. Two-sample nonparametric test for comparing mean time to failure functions in age replacement. Journal of Statistical Planning and Inference, 212(??):34–44, May 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820301075.

Baek:2020:ABL

[BKP20] Changryong Baek, Stefanos Kechagias, and Vladas Pipiras. Asymptotics of bivariate local Whittle estimators with applications to fractal connectivity. Journal of Statistical Planning and Inference, 205(??):245–268, March 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375819300734.

Bachoc:2020:FDA

[BL20] F. Bachoc and A. Lagnoux. Fixed-domain asymptotic properties of maximum composite likelihood estimators for Gaussian processes. Journal of Statistical Planning and Inference, 209(??):62– 75, December 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375820300239.

Bohringer:2022:EMC

[BL22] Stefan Böhringer and Dietmar Lohmann. Exact model comparisons in the plausibility framework. Journal of Statistical Planning and Inference, 217(??):224–240, March 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837582100080X.

Brairi:2020:TDV

[BM20] Houssem Brairi and Tarek Medkour. Testing discrete-valued time series for whiteness. Journal of Statistical Planning and Inference, 206(??):43–56, May 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375819300795.

Banerjee:2021:OEE

[BM21] Samprit Banerjee and Stefano Monni. An orthogonally equivariant estimator of the covariance matrix in high dimensions and for small sample sizes. Journal of Statistical Planning and Inference, 213 (??):16–32, July 2021. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375820301142.

Bhattacharya:2022:GPI

[BM22] Indrabati Bhattacharya and Ryan Martin. Gibbs posterior inference on multivariate quantiles. Journal of Statistical Planning and Inference, 218(??):106–121, May 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821001051.

Beran:2023:SFM

[BNF^{+}23] Jan Beran, Jeremy Näscher, Franziska Farquharson, Max Kustermann, Hans-Joachim Kabitz, and Stephan Walterspacher. On seasonal functional modeling under strong dependence, with applications to mechanically ventilated breathing activity. Journal of Statistical Planning and Inference, 222(??):38–65, January 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000453.

Britos:2020:APB

[BOAB20] Grisel M. Britos, Silvia M. Ojeda, Laura A. Rodríguez Astrain, and Oscar H. Bustos. Asymptotic properties of BMMestimator in bidimensional autoregressive processes. Journal of Statistical Planning and Inference, 209(??):208–228, December 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300392.

Benke:2021:NUF

[BP21] János Marcell Benke and Gyula Pap. Nearly unstable family of stochastic processes given by stochastic differential equations with time delay. Journal of Statistical Planning and Inference, 211(??): 1–11, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375820300653.

Bailey:2021:USL

[BS21a] R. A. Bailey and Leonard H. Soicher. Uniform semi-Latin squares and their pairwise-variance aberrations. Journal of Statistical Planning and Inference, 213(??):282–291, July 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820301373.

Bentarzi:2021:EEP

[BS21b] Mohamed Bentarzi and Mohamed Sadoun. Efficient estimation in periodic INARp model: Nonparametric innovation distributions case. Journal of Statistical Planning and Inference, 211(??):340– 361, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375820300872.

Branson:2021:RRE

[BS21c] Zach Branson and Stephane Shao. Ridge rerandomization: an experimental design strategy in the presence of covariate collinearity. Journal of Statistical Planning and Inference, 211(??):287– 314, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375820300847.

Banerjee:2022:GSD

[BS22] Sayantan Banerjee and Weining Shen. Graph signal denoising using t-shrinkage priors. Journal of Statistical Planning and Inference, 219(??):279–305, July 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837582100135X.

Buonaguidi:2020:DPP

[Buo20] B. Buonaguidi. The disorder problem for purely jump Lévy processes with completely monotone jumps. Journal of Statistical Planning and Inference, 205(??):203–218, March 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375819300709.

Cavicchioli:2022:GFT

[Cav22] Maddalena Cavicchioli. Goodness-of-fit tests for Markov Switching VAR models using spectral analysis. Journal of Statistical Planning and Inference, 219(??):189–203, July 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821001324.

Chatterjee:2020:RMG

[CB20] A. Chatterjee and T. Bandyopadhyay. Regression models for group testing: Identifiability and asymptotics. Journal of Statistical Planning and Inference, 204(??):141–152, January 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818301423.

Chatterjee:2024:IRM

[CBB24] Arindam Chatterjee, Tathagata Bandyopadhyay, and Ayoushman Bhattacharya. Inference on regression model with misclassified binary response. Journal of Statistical Planning and Inference, 231(??):Article 106121, July 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375823000903.

Cordoba:2020:RGM

[CBL20] Irene Córdoba, Concha Bielza, and Pedro Larrañaga. A review of Gaussian Markov models for conditional independence. Journal of Statistical Planning and Inference, 206(??):127–144, May 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819300916.

Chagny:2022:ANE

[CCHR22] Gaëlle Chagny, Antoine Channarond, Van Hà Hoang, and Angelina Roche. Adaptive nonparametric estimation of a component density in a two-class mixture model. Journal of Statistical Planning and Inference, 216(??):51–69, January 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000495.

Cuevas:2020:NDU

[CF20] Antonio Cuevas and Ricardo Fraiman. Nonparametric detection for univariate and functional data. Journal of Statistical Planning and Inference, 209(??):12–26, December 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820300203.

Cardot:2024:HTP

[CF24] Hervé Cardot and Cindy Frascolla. Hypothesis testing for panels of semi-Markov processes with parametric sojourn time distributions. Journal of Statistical Planning and Inference, 228(??): 59–79, January 2024. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375823000459.

Cholaquidis:2023:UCE

[CFM23] Alejandro Cholaquidis, Ricardo Fraiman, and Leonardo Moreno. Universally consistent estimation of the reach. Journal of Statistical Planning and Inference, 225(??):110–120, July 2023. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837582200101X.

Chauvet:2022:AEC

[CG22] Guillaume Chauvet and Camelia Goga. Asymptotic efficiency of the calibration estimator in a high-dimensional data setting. Journal of Statistical Planning and Inference, 217(??):177–187, March 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000781.

Chen:2024:FSC

[Che24] Li-Pang Chen. Feature screening via concordance indices for left-truncated and right-censored survival data. Journal of Statistical Planning and Inference, 232(??):Article 106153, September 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375824000107.

Chu:2023:LLR

[Chu23] Ba Chu. Local linear regression with nonparametrically generated covariates for weakly dependent data. Journal of Statistical Planning and Inference, 225(??):89–109, July 2023. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375822001008.

Clertant:2024:EEP

[CIO24] Matthieu Clertant, Alexia Iasonos, and John O’Quigley. Evaluation of early phase dose finding algorithms in heterogeneous populations. Journal of Statistical Planning and Inference, 231(??):Article 106080, July 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375823000356.

Cai:2021:OSE

[CKSW21] Tony Cai, Donggyu Kim, Xinyu Song, and Yazhen Wang. Optimal sparse eigenspace and low-rank density matrix estimation for quantum systems. Journal of Statistical Planning and Inference, 213 (??):50–71, July 2021. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375820301166.

Clairon:2021:RMP

[Cla21] Quentin Clairon. A regularization method for the parameter estimation problem in ordinary differential equations via discrete optimal control theory. Journal of Statistical Planning and Inference, 210(??):1–19, January 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820300458.

Chen:2023:SSM

[CLPX23] Shuaiyu Chen, Yan Li, Xiaolong Pu, and Dongdong Xiang. A stable sequential multiple test for Koopman–Darmois family. Journal of Statistical Planning and Inference, 226(??):39–62, September 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837582300006X.

Chen:2023:MEA

[CLT23] Yong Chen, Ying Li, and Li Tian. Moment estimator for an AR(1) model driven by a long memory Gaussian noise. Journal of Statistical Planning and Inference, 222(??):94–107, January 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000507.

Comte:2020:BSW

[CM20] Fabienne Comte and Nicolas Marie. Bandwidth selection for the wolverton-Wagner estimator. Journal of Statistical Planning and Inference, 207(??):198–214, July 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375819301223.

Chen:2023:PUS

[CM23] Xin Chen and Mary C. Meyer. Penalized unimodal spline density estimation with application to M-estimation. Journal of Statistical Planning and Inference, 224(??):84–97, May 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000933.

Columbu:2021:HSR

[CMMD21] Silvia Columbu, Valentina Mameli, Monica Musio, and Philip Dawid. The Hyvärinen scoring rule in Gaussian linear time series models. Journal of Statistical Planning and Inference, 212(??): 126–140, May 2021. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837582030104X.

Chen:2020:KDR

[CMZ20] Xin Chen, Xuejun Ma, and Wang Zhou. Kernel density regression. Journal of Statistical Planning and Inference, 205(??):318– 329, March 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375818303264.

Chen:2022:VSB

[CQY22] Baojiang Chen, Jing Qin, and Ao Yuan. Variable selection in the Box–Cox power transformation model. Journal of Statistical Planning and Inference, 216(??):15–28, January 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000483.

Chen:2020:BHP

[CS20] Xiongzhi Chen and Sanat K. Sarkar. On Benjamini–Hochberg procedure applied to mid p-values. Journal of Statistical Planning and Inference, 205(??):34–45, March 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375819300540.

Cai:2020:GVS

[CSL20] Kaida Cai, Hua Shen, and Xuewen Lu. Group variable selection in the Andersen–Gill model for recurrent event data. Journal of Statistical Planning and Inference, 207(??):99–112, July 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819301144.

Cao:2021:STM

[CSP21] Mingxiang Cao, Peng Sun, and Junyong Park. A simultaneous test of mean vector and covariance matrix in high-dimensional settings. Journal of Statistical Planning and Inference, 212(??): 141–152, May 2021. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375820301051.

Chen:2022:USO

[CT22a] E. Chen and Yu Tang. Uniformity of saturated orthogonal arrays. Journal of Statistical Planning and Inference, 216(??):174– 181, January 2022. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375821000677.

Chen:2022:UND

[CT22b] Guanzhou Chen and Boxin Tang. Using nonregular designs to generate space-filling designs. Journal of Statistical Planning and Inference, 221(??):201–211, December 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375822000374.

Chen:2020:SWC

[CWS20] Pingyan Chen, Luliang Wen, and Soo Hak Sung. Strong and weak consistency of least squares estimators in simple linear EV regression models. Journal of Statistical Planning and Inference, 205(??): 64–73, March 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375819300655.

Cheng:2020:IBO

[CWY20] Qianshun Cheng, HaiYing Wang, and Min Yang. Informationbased optimal subdata selection for big data logistic regression. Journal of Statistical Planning and Inference, 209(??):112–122, December 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375820300331.

DeOliveira:2021:BDG

[DCDS21] Melaine C. De Oliveira, Luis M. Castro, Dipak K. Dey, and Debajyoti Sinha. Bregman divergence to generalize Bayesian influence measures for data analysis. Journal of Statistical Planning and Inference, 213(??):222–232, July 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820301257.

Detmer:2020:NCS

[DCS20] Felicitas J. Detmer, Juan Cebral, and Martin Slawski. A note on coding and standardization of categorical variables in (sparse) group lasso regression. Journal of Statistical Planning and Inference, 206(??):1–11, May 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375819300758.

Diaz-Garcia:2021:MVB

[DGCL21] José A. Díaz-García and Francisco J. Caro-Lopera. Matrix variate Birnbaum–Saunders distribution under elliptical models. Journal of Statistical Planning and Inference, 210(??):100–113, January 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300562.

Diaz-Garcia:2022:MDU

[DGCL22] José A. Díaz-García and Francisco J. Caro-Lopera. Multimatricvariate distribution under elliptical models. Journal of Statistical Planning and Inference, 216(??):109–117, January 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000628.

Domagni:2021:NSF

[DHS21] Kouakou Francois Domagni, A. S. Hedayat, and Bikas Kumar Sinha. On the nature of saturated 2^{k}-factorial designs for unbiased estimation of non-negligible parameters. Journal of Statistical Planning and Inference, 211(??):33–40, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300641.

Dao:2022:VEC

[DJPZ22] Cecilia Dao, Jiming Jiang, Debashis Paul, and Hongyu Zhao. Variance estimation and confidence intervals from genome-wide association studies through high-dimensional misspecified mixed model analysis. Journal of Statistical Planning and Inference, 220(??):15– 23, September 2022. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375822000039.

Diop:2021:PAG

[DK21] Mamadou Lamine Diop and William Kengne. Piecewise autoregression for general integer-valued time series. Journal of Statistical Planning and Inference, 211(??):271–286, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300859.

Dukes:2023:GDM

[DLZ23] Peter Dukes, Meixin Liu, and Julie Zhou. Graph decomposition methods for variance balanced block designs with correlated errors. Journal of Statistical Planning and Inference, 222(??):252– 260, January 2023. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375822000660.

Dasgupta:2022:ODS

[DMK22] Subhadra Dasgupta, Siuli Mukhopadhyay, and Jonathan Keith. Optimal designs for some bivariate cokriging models. Journal of Statistical Planning and Inference, 221(??):9–28, December 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000167.

Dong:2021:BRL

[Don21] Yuexiao Dong. A brief review of linear sufficient dimension reduction through optimization. Journal of Statistical Planning and Inference, 211(??):154–161, March 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837582030077X.

Durand:2024:HSV

[DR24] Amaury Durand and François Roueff. Hilbert space-valued fractionally integrated autoregressive moving average processes with long memory operators. Journal of Statistical Planning and Inference, 231(??):Article 106146, July 2024. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837582400003X.

Das:2020:DCE

[DS20a] Ashish Das and Rakhi Singh. Discrete choice experiments — a unified approach. Journal of Statistical Planning and Inference, 205(??):193–202, March 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375819300692.

Delevoye:2020:CHT

[DS20b] Angèle Delevoye and Fredrik Sävje. Consistency of the Horvitz– Thompson estimator under general sampling and experimental designs. Journal of Statistical Planning and Inference, 207(??): 190–197, July 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375819301211.

Benaduce:2023:SDM

[dSCBP23] Helen da Silva Costa Benaduce and Guilherme Pumi. SYMARFIMA: a dynamical model for conditionally symmetric time series with long range dependence mean structure. Journal of Statistical Planning and Inference, 225(??):71–88, July 2023. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375822000969.

Dumitrescu:2021:ARE

[DSK21] Laura Dumitrescu and Ioana Schiopu-Kratina. Asymptotic results with estimating equations for time-evolving clustered data. Journal of Statistical Planning and Inference, 214(??):41–61, September 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000070.

Dong:2023:SRS

[DSP23] Yuexiao Dong, Abdul-Nasah Soale, and Michael D. Power. A selective review of sufficient dimension reduction for multivariate response regression. Journal of Statistical Planning and Inference, 226(??):63–70, September 2023. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375823000150.

Dupuis:2022:POS

[DT22] Xavier Dupuis and Patrick J. C. Tardivel. Proximal operator for the sorted ℓ_{1} norm: Application to testing procedures based on SLOPE. Journal of Statistical Planning and Inference, 221(??): 1–8, December 2022. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375822000179.

Canditiis:2024:ALE

[DT24] Daniela De Canditiis and Giovanni Luca Torrisi. On the adaptive lasso estimator of AR(p) time series with applications to INAR(p) and Hawkes processes. Journal of Statistical Planning and Inference, 231(??):Article 106145, July 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375824000028.

Dussap:2021:AMD

[Dus21] Florian Dussap. Anisotropic multivariate deconvolution using projection on the Laguerre basis. Journal of Statistical Planning and Inference, 215(??):23–46, December 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000215.

Deng:2024:MFM

[DZ24] Linsui Deng and Yilin Zhang. Model-free, monotone invariant and computationally efficient feature screening with data-adaptive threshold. Journal of Statistical Planning and Inference, 228(??): 23–33, January 2024. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375823000472.

Machkouri:2020:NWK

[EFR20] M. El Machkouri, X. Fan, and L. Reding. On the Nadaraya– Watson kernel regression estimator for irregularly spaced spatial data. Journal of Statistical Planning and Inference, 205(??):92– 114, March 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375818301435.

Eck:2020:CEM

[EGC20] Daniel J. Eck, Charles J. Geyer, and R. Dennis Cook. Combining envelope methodology and aster models for variance reduction in life history analyses. Journal of Statistical Planning and Inference, 205(??):283–292, March 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375818302374.

ElBarmi:2021:EDF

[EMM21] Hammou El Barmi, Ganesh Malla, and Hari Mukerjee. Estimation of a distribution function with increasing failure rate average. Journal of Statistical Planning and Inference, 213(??): 179–192, July 2021. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375820301026.

Fontana:2020:OBC

[FC20] Roberto Fontana and Francesca Romana Crucinio. Orbit-based conditional tests. A link between permutations and Markov bases. Journal of Statistical Planning and Inference, 205(??): 23–33, March 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375819300539.

Fernandez-Duran:2024:RMC

[FDGD24] Juan José Fernández-Durán and María Mercedes Gregorio-Domínguez. Regression models for circular data based on nonnegative trigonometric sums. Journal of Statistical Planning and Inference, 230(??):Article 106114, May 2024. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375823000836.

Florent:2023:SNO

[FE23] Kouakou Kouamé Florent and Yodé Armel Fabrice Evrard. A supnorm oracle inequality for a partially linear regression model. Journal of Statistical Planning and Inference, 222(??):132–148, January 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000611.

Fuchs:2020:ABV

[FHBD20] Mathias Fuchs, Roman Hornung, Anne-Laure Boulesteix, and Riccardo De Bin. On the asymptotic behaviour of the variance estimator of a U-statistic. Journal of Statistical Planning and Inference, 209(??):101–111, December 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837582030032X.

Fu:2024:RPE

[FHL24] Liya Fu, Shuwen Hu, and Jiaqi Li. Robust penalized empirical likelihood in high dimensional longitudinal data analysis. Journal of Statistical Planning and Inference, 228(??):11–22, January 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823000435.

Fasen-Hartmann:2022:NES

[FHM22] Vicky Fasen-Hartmann and Celeste Mayer. A note on estimation of α-stable CARMA processes sampled at low frequencies. Journal of Statistical Planning and Inference, 219(??):250–265, July 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821001361.

Freise:2020:ODT

[FHS20] Fritjof Freise, Heinz Holling, and Rainer Schwabe. Optimal designs for two-level main effects models on a restricted design region. Journal of Statistical Planning and Inference, 204(??):45– 54, January 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375819300370.

Freise:2023:ODE

[FHS23] Fritjof Freise, Heinz Holling, and Rainer Schwabe. Optimal design for estimating the mean ability over time in repeated item response testing. Journal of Statistical Planning and Inference, 225(??): 266–282, July 2023. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375823000034.

Fang:2024:DEA

[FLL^{+}24] Zhiwen Fang, Gen Li, Wendong Li, Xiaolong Pu, and Dongdong Xiang. Distributed eQTL analysis with auxiliary information. Journal of Statistical Planning and Inference, 228(??):34– 45, January 2024. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375823000447.

Flournoy:2021:EAM

[FMT21] Nancy Flournoy, Caterina May, and Chiara Tommasi. The effects of adaptation on maximum likelihood inference for nonlinear models with normal errors. Journal of Statistical Planning and Inference, 214(??):139–150, September 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000185.

Frolov:2022:LDC

[Fro22] Andrei N. Frolov. On large deviations for combinatorial sums. Journal of Statistical Planning and Inference, 217(??):24–32, March 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000690.

Fontana:2023:EBD

[FS23] R. Fontana and P. Semeraro. Exchangeable Bernoulli distributions: High dimensional simulation, estimation, and testing.

Journal of Statistical Planning and Inference, 225(??):52–70, July 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000994.

Feng:2021:EFS

[FTHL21] Sanying Feng, Ping Tian, Yuping Hu, and Gaorong Li. Estimation in functional single-index varying coefficient model. Journal of Statistical Planning and Inference, 214(??):62–75, September 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000045.

Fang:2022:RMV

[FY22] Junhan Fang and Grace Y. Yi. Regularized matrix-variate logistic regression with response subject to misclassification. Journal of Statistical Planning and Inference, 217(??):106–121, March 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000689.

Fu:2024:MCD

[FZ24] Lianyan Fu and Luyang Zhang. Measures of conditional dependence for nonlinearity, asymmetry and beyond. Journal of Statistical Planning and Inference, 232(??):Article 106165, September 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375824000223.

Garcia-Angulo:2023:OFS

[GAC23] Andrea C. Garcia-Angulo and Gerda Claeskens. Optimal finite sample post-selection confidence distributions in generalized linear models. Journal of Statistical Planning and Inference, 222(??):66– 77, January 2023. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375822000489.

Gelein:2020:PDF

[GC20] Brigitte Gelein and Guillaume Chauvet. Preserving the distribution function in surveys in case of imputation for zero inflated data. Journal of Statistical Planning and Inference, 206 (??):84–99, May 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375819300886.

Gong:2021:SAF

[GC21] Yuan Gong and Zehua Chen. A sequential approach to feature selection in high-dimensional additive models. Journal of Statistical Planning and Inference, 215(??):289–298, December 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000434.

Ghosh:2020:GBE

[GD20]  Sayan Ghosh and Ori Davidov.  Graph-based estimators

for paired comparison data. Journal of Statistical Planning and Inference, 209(??):1–11, December 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820300197.

Grasshoff:2020:ORD

[GHRS20] Ulrike Graßhoff, Heinz Holling, Frank Röttger, and Rainer Schwabe. Optimality regions for designs in multiple linear regression models with correlated random coefficients. Journal of Statistical Planning and Inference, 209(??):267–279, December 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300422.

Gaucher:2021:MLE

[GK21] Solenne Gaucher and Olga Klopp. Maximum likelihood estimation of sparse networks with missing observations. Journal of Statistical Planning and Inference, 215(??):299–329, December 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000422.

Geng:2022:WEM

[GK22] Pei Geng and Hira L. Koul. Weighted empirical minimum distance estimators in linear errors-in-variables regression models. Journal of Statistical Planning and Inference, 219(??):147–174, July 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821001312.

Gijbels:2022:CSP

[GKO22] Irène Gijbels, Vojtech Kika, and Marek Omelka. Choice of smoothing parameter in multivariate copula-based tail coefficients. Journal of Statistical Planning and Inference, 221(??):136–153, December 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000325.

Guo:2020:QLE

[GL20] Feifei Guo and Shiqing Ling. Quasi-likelihood estimation of structure-changed threshold double autoregressive models. Journal of Statistical Planning and Inference, 205(??):138–155, March 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819300679.

Gadrich:2021:TWO

[GM21] Tamar Gadrich and Yariv N. Marmor. Two-way ORDANOVA: Analyzing ordinal variation in a cross-balanced design. Journal of Statistical Planning and Inference, 215(??):330–343, December 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000446.

Geng:2020:EPL

[GMFY20] Xin Geng, Carlos Martins-Filho, and Feng Yao. Estimation of a partially linear additive model with generated covariates. Journal of Statistical Planning and Inference, 208(??):94–118, September 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300173.

Gorajski:2024:THI

[Gór24] Mariusz Górajski. Testing higher and infinite degrees of stochastic dominance for small samples: a Bayesian approach. Journal of Statistical Planning and Inference, 230(??):Article 106102, May 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837582300071X.

Grant:2020:CTT

[GPG20a] Sheridan Grant, Michael Perlman, and Darren Grant. Corrigendum to “Targeted testing for bias in order assignment, with an application to Texas election ballots” [J. Statist. Plann. Inference (2020) 12–28]. Journal of Statistical Planning and Inference, 207 (??):344, July 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375819301120. See [GPG20b].

Grant:2020:TTB

[GPG20b] Sheridan Grant, Michael D. Perlman, and Darren Grant. Targeted testing for bias in order assignment, with an application to Texas election ballots. Journal of Statistical Planning and Inference, 206 (??):12–28, May 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375818300867. See corrigendum

[GPG20a].

Ghosh:2023:ASO

[GPP23] Subir Ghosh, Luyao Peng, and Hans-Peter Piepho. An algorithm for searching optimal variance component estimators in linear mixed models. Journal of Statistical Planning and Inference, 227(??):34–56, December 2023. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375823000198.

Guha:2024:RAB

[GR24] Sharmistha Guha and Jerome P. Reiter. Regression-assisted Bayesian record linkage for causal inference in observational studies with covariates spread over two files. Journal of Statistical Planning and Inference, 229(??):Article 106090, March 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823000514.

Grossmann:2020:MBE

[Gro20] Heiko Großmann. On the meaning of block effects in paired comparison choice experiments and a relationship with blocked 2^{K} main effects plans. Journal of Statistical Planning and Inference, 209(??):76–84, December 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375820300240.

Guo:2020:ACF

[GS20] Wenge Guo and Sanat Sarkar. Adaptive controls of FWER and FDR under block dependence. Journal of Statistical Planning and Inference, 208(??):13–24, September 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375819301181.

Girardin:2024:TCS

[GS24] V. Girardin and R. Senoussi. Time changes and stationarity issues for extended scalar autoregressive models. Journal of Statistical Planning and Inference, 230(??):Article 106112, May 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823000812.

Gao:2020:MOD

[GZ20] Lucy L. Gao and Julie Zhou. Minimax D-optimal designs for multivariate regression models with multi-factors. Journal of Statistical Planning and Inference, 209(??):160–173, December 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300367.

Garbuzov:2023:CRP

[GZ23] Boris Garbuzov and Zhou Zhou. Convergence rate of plugin estimates for functional parameters with applications to locally-stationary time-series. Journal of Statistical Planning and Inference, 223(??):53–62, March 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375822000799.

Gao:2020:SMR

[GZZ20] Lei Gao, Hongjian Zhu, and Lanju Zhang. Sequential monitoring of response-adaptive randomized clinical trials with sample size reestimation. Journal of Statistical Planning and Inference, 205(??): 129–137, March 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375818301873.

Haines:2022:AOD

[Hai22] Linda M. Haines. Approximate I-optimal designs for polynomial models over the unit ball. Journal of Statistical Planning and Inference, 220(??):71–81, September 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375822000155.

Hoyos-Arguelles:2020:BSA

[HANB20] Ricardo Hoyos-Argüelles and Luis Nieto-Barajas. A Bayesian semiparametric Archimedean copula. Journal of Statistical Planning and Inference, 206(??):298–311, May 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375819301004.

Hashimoto:2021:RPD

[Has21] Shintaro Hashimoto. Reference priors via α-divergence for a certain non-regular model in the presence of a nuisance parameter. Journal of Statistical Planning and Inference, 213(??):162– 178, July 2021. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375820301221.

He:2021:AOS

[HB21] Xinrui He and Jay Bartroff. Asymptotically optimal sequential FDR and pFDR control with (or without) prior information on the number of signals. Journal of Statistical Planning and Inference, 210(??):87–99, January 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820300604.

Harrar:2023:NMC

[HC23] Solomon W. Harrar and Yue Cui. Nonparametric methods for clustered data in pre-post intervention design. Journal of Statistical Planning and Inference, 222(??):1–21, January 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000477.

He:2023:HDV

[HCX23] Daojiang He, Jinjiao Cheng, and Kai Xu. High-dimensional variable screening through kernel-based conditional mean dependence. Journal of Statistical Planning and Inference, 224(??): 27–41, May 2023. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375822000908.

Hu:2024:OMA

[HCZG24] Guozhi Hu, Weihu Cheng, Jie Zeng, and Ruijie Guan. Optimal model averaging for semiparametric partially linear models with measurement errors. Journal of Statistical Planning and Inference, 230(??):Article 106101, May 2024. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375823000708.

Hu:2020:SIL

[HDLF20] Jianhua Hu, Hao Ding, Liangyuan Liu, and Jingyan Feng. Statistical inference of locally stationary functional coefficient models. Journal of Statistical Planning and Inference, 209(??):27– 43, December 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375820300215.

Heller:2023:MNR

[Hel23] Glenn Heller. A modified net reclassification improvement statistic. Journal of Statistical Planning and Inference, 227(??):18– 33, December 2023. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375823000186.

Hsu:2021:ODE

[HH21] Hsiang-Ling Hsu and Mong-Na Lo Huang. D-optimal designs for estimation of parameters in a simplex dispersion model with proportional data. Journal of Statistical Planning and Inference, 215(??):193–207, December 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000355.

He:2024:FZI

[HH24] Qing He and Hsin-Hsiung Huang. A framework of zero-inflated Bayesian negative binomial regression models for spatiotemporal data. Journal of Statistical Planning and Inference, 229(??):Article 106098, March 2024. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375823000678.

Huang:2020:ODB

[HHL20] Shih-Hao Huang, Mong-Na Lo Huang, and Cheng-Wei Lin. Optimal designs for binary response models with multiple nonnegative variables. Journal of Statistical Planning and Inference, 206 (??):75–83, May 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375819300813.

Hoberman:2022:PEM

[HI22] Steven Hoberman and Anastasia Ivanova. The properties of entropy as a measure of randomness in a clinical trial. Journal of Statistical Planning and Inference, 216(??):182–193, January 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000604.

Hill:2024:SVT

[Hil24] Jonathan B. Hill. A smoothed p-value test when there is a nuisance parameter under the alternative. Journal of Statistical Planning and Inference, 229(??):Article 106096, March 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823000630.

Hamura:2022:BPD

[HK22] Yasuyuki Hamura and Tatsuya Kubokawa. Bayesian predictive density estimation for a chi-squared model using information from a normal observation with unknown mean and variance. Journal of Statistical Planning and Inference, 217(??):33– 51, March 2022. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375821000719.

Hu:2022:EEL

[HLDB22] Jie Hu, Wei Liang, Hongsheng Dai, and Yanchun Bao. Efficient empirical likelihood inference for recovery rate of COVID19 under double-censoring. Journal of Statistical Planning and Inference, 221(??):172–187, December 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375822000350.

Han:2024:RIS

[HLLW24] Miao Han, Yuanyuan Lin, Wenxin Liu, and Zhanfeng Wang. Robust inference for subgroup analysis with general transformation models. Journal of Statistical Planning and Inference, 229(??):Article 106100, March 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375823000691.

Huang:2023:WBT

[HLP23] Bingyao Huang, Yanyan Liu, and Liuhua Peng. Weighted bootstrap for two-sample U-statistics. Journal of Statistical Planning and Inference, 226(??):86–99, September 2023. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375823000162.

Han:2020:MLF

[HLYZ20] Xiaoxue Han, Min-Qian Liu, Jian-Feng Yang, and Shengli Zhao. Mixed 2- and 2r-level fractional factorial split-plot designs with clear effects. Journal of Statistical Planning and Inference, 204(??): 206–216, January 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375819300527.

He:2024:CHD

[HS24a] Xu He and Fasheng Sun. Construction of high-dimensional high-separation distance designs. Journal of Statistical Planning and Inference, 232(??):Article 106150, September 2024. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375824000077.

Hewage:2024:JEL

[HS24b] Sameera Hewage and Yongli Sang. Jackknife empirical likelihood confidence intervals for the categorical Gini correlation. Journal of Statistical Planning and Inference, 231(??):Article 106123, July 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823000927.

Huang:2021:CCO

[Hua21] Hengzhen Huang. Construction of component orthogonal arrays with any number of components. Journal of Statistical Planning and Inference, 213(??):72–79, July 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820301178.

Hu:2024:CML

[HWS24] Bo Hu, Dongying Wang, and Fasheng Sun. Construction of mixed-level screening designs using Hadamard matrices. Journal of Statistical Planning and Inference, 231(??):Article 106131, July 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823001003.

Hu:2020:SNP

[HXZZ20] Yuping Hu, Liugen Xue, Jing Zhao, and Liying Zhang. Skewnormal partial functional linear model and homogeneity test. Journal of Statistical Planning and Inference, 204(??):116–127, January 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818301903.

Hao:2020:SPL

[HY20] Shuai Hao and Min Yang. Support point of locally optimal designs for multinomial logistic regression models. Journal of Statistical Planning and Inference, 209(??):144–159, December 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300355.

He:2024:ODC

[HYD24] Lei He, Rong-Xian Yue, and Andrew Du. Optimal designs for comparing curves in regression models with asymmetric errors. Journal of Statistical Planning and Inference, 228(??):46–58, January 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823000460.

Ito:2021:CEB

[IK21] Tsubasa Ito and Tatsuya Kubokawa. Corrected empirical Bayes confidence region in a multivariate Fay–Herriot model. Journal of Statistical Planning and Inference, 211(??):12–32, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818301617.

Jazi:2024:ECM

[Jaz24] Omidali Aghababaei Jazi. Estimation for the Cox model with biased sampling data via risk set sampling. Journal of Statistical Planning and Inference, 228(??):1–10, January 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823000423.

Jeong:2022:PCG

[Jeo22] Seonghyun Jeong. Posterior contraction in group sparse logit models for categorical responses. Journal of Statistical Planning and Inference, 219(??):266–278, July 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375822000015.

Jeong:2023:NMO

[Jeo23] Minsoo Jeong. A numerical method to obtain exact confidence intervals for likelihood-based parameter estimators. Journal of Statistical Planning and Inference, 226(??):20–29, September 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822001136.

Jiang:2021:LSS

[JHH21] Dandan Jiang, Zhiqiang Hou, and Jiang Hu. The limits of the sample spiked eigenvalues for a high-dimensional generalized Fisher matrix and its applications. Journal of Statistical Planning and Inference, 215(??):208–217, December 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000379.

Jiang:2020:MPC

[JK20] Peiyun Jiang and Eiji Kurozumi. Monitoring parameter changes in models with a trend. Journal of Statistical Planning and Inference, 207(??):288–319, July 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820300021.

Johansson:2022:RCS

[JS22] Per Johansson and Mårten Schultzberg. Rerandomization: a complement or substitute for stratification in randomized experiments? Journal of Statistical Planning and Inference, 218(??): 43–58, May 2022. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837582100094X.

Jauslin:2023:EAS

[JT23] Raphaël Jauslin and Yves Tillé. An efficient approach for statistical matching of survey data through calibration, optimal transport and balanced sampling. Journal of Statistical Planning and Inference, 225(??):121–131, July 2023. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375822001100.

Jiang:2022:TSS

[JX22] Yuanyuan Jiang and Xingzhong Xu. Testing the skewness of skew-normal distribution by Bayes factors. Journal of Statistical Planning and Inference, 220(??):24–48, September 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000143.

Kakizawa:2020:MNC

[Kak20]  Yoshihide Kakizawa.  Multivariate non-central Birnbaum–

Saunders kernel density estimator for nonnegative data. Journal of Statistical Planning and Inference, 209(??):187–207, December 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300380.

Kalemkerian:2023:MPE

[Kal23] Juan Kalemkerian. Modelling and parameter estimation for discretely observed fractional iterated Ornstein–Uhlenbeck processes. Journal of Statistical Planning and Inference, 225(??):29–51, July 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000982.

Krishnamoorthy:2024:CIR

[KC24] K. Krishnamoorthy and Saptarshi Chakraberty. Confidence intervals for a ratio of percentiles of location-scale distributions. Journal of Statistical Planning and Inference, 229(??):Article 106089, March 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823000502.

Kwan:2023:AAI

[KCD23] Tsz-Kit Jeffrey Kwan, Feng Chen, and William T. M. Dunsmuir. Alternative asymptotic inference theory for a nonstationary Hawkes process. Journal of Statistical Planning and Inference, 227(??):75–90, December 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375823000307.

Kim:2021:SIP

[KCH21] Kun Ho Kim, Shih-Kang Chao, and Wolfgang K. Härdle. Simultaneous inference of the partially linear model with a multivariate unknown function. Journal of Statistical Planning and Inference, 213 (??):93–105, July 2021. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375820301208.

Kuwada:2020:LBN

[KH20] Masahide Kuwada and Yoshifumi Hyodo. On a lower bound for the number of assemblies in fractional 2^{m} factorial designs of resolution 2ℓ. Journal of Statistical Planning and Inference, 207(??): 181–189, July 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837581930120X.

Khmaladze:2021:THT

[Khm21] Estate V. Khmaladze. Testing hypothesis on transition distributions of a Markov sequence. Journal of Statistical Planning and Inference, 215(??):72–84, December 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000252.

Kim:2020:MGF

[Kim20] Ilmun Kim. Multinomial goodness-of-fit based on U-statistics: High-dimensional asymptotic and minimax optimality. Journal of Statistical Planning and Inference, 205(??):74–91, March 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819300667.

Kariya:2024:MFR

[KKH24] Takeaki Kariya, Hiroshi Kurata, and Takaki Hayashi. A modelling framework for regression with collinearity. Journal of Statistical Planning and Inference, 228(??):95–115, January 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823000484.

Kapelner:2022:ORD

[KKSA22] Adam Kapelner, Abba M. Krieger, Michael Sklar, and David Azriel. Optimal rerandomization designs via a criterion that provides insurance against failed experiments. Journal of Statistical Planning and Inference, 219(??):63–84, July 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821001130.

Kong:2023:RCS

[KLCZ23] Qingxun Kong, Zhiming Li, Tao Chen, and Ge Zhang. Results on constructing sn-m regular designs with general minimum lower-order confounding. Journal of Statistical Planning and Inference, 222(??):122–131, January 2023. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375822000490.

Kohler:2023:ERF

[KLR23] Michael Kohler, Sophie Langer, and Ulrich Reif. Estimation of a regression function on a manifold by fully connected deep neural networks. Journal of Statistical Planning and Inference, 222(??): 160–181, January 2023. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375822000465.

Kateri:2024:STM

[KN24] Maria Kateri and Nikolay I. Nikolov. Scale tests for a multilevel step-stress model with exponential lifetimes under Type-II censoring. Journal of Statistical Planning and Inference, 232(??):Article 106152, September 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375824000090.

Kevei:2022:LLN

[KOV22] Péter Kevei, Lillian Oluoch, and László Viharos. Limit laws for the norms of extremal samples. Journal of Statistical Planning and Inference, 216(??):151–173, January 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000665.

Kuba:2024:CGT

[KP24] Markus Kuba and Alois Panholzer. On card guessing with two types of cards. Journal of Statistical Planning and Inference, 232(??):Article 106160, September 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837582400017X.

Kolkiewicz:2021:PPB

[KRX21] Adam Kolkiewicz, Gregory Rice, and Yijun Xie. Projection pursuit based tests of normality with functional data. Journal of Statistical Planning and Inference, 211(??):326–339, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300835.

Kirch:2022:ADT

[KS22] Claudia Kirch and Christina Stoehr. Asymptotic delay times of sequential tests based on U-statistics for early and late change points. Journal of Statistical Planning and Inference, 221(??):114– 135, December 2022. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375822000313.

Kaino:2021:PEP

[KU21] Yusuke Kaino and Masayuki Uchida. Parametric estimation for a parabolic linear SPDE model based on discrete observations. Journal of Statistical Planning and Inference, 211(??):190–220, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300628.

Kawai:2022:ATM

[KU22] Tetsuya Kawai and Masayuki Uchida. Adaptive testing method for ergodic diffusion processes based on high frequency data. Journal of Statistical Planning and Inference, 217(??):241–278, March 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000914.

Kusano:2024:SIF

[KU24] Shogo Kusano and Masayuki Uchida. Statistical inference in factor analysis for diffusion processes from discrete observations. Journal of Statistical Planning and Inference, 229(??):Article 106095, March 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823000563.

Kalogridis:2021:TPS

[KV21] Ioannis Kalogridis and Stefan Van Aelst. M-type penalized splines with auxiliary scale estimation. Journal of Statistical Planning and Inference, 212(??):97–113, May 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820301087.

Karmakar:2023:NCI

[KVH^{+}23] Sayantani Karmakar, Cini Varghese, Md. Ashraful Haque, Seema Jaggi, Mohd Harun, and Eldho Varghese. A note on the construction of incomplete row-column designs: an algorithmic approach. Journal of Statistical Planning and Inference, 222(??): 108–121, January 2023. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375822000519.

Kleijn:2023:AUQ

[KvW23] B. J. K. Kleijn and J. van Waaij. Asymptotic uncertainty quantification for communities in sparse planted bi-section models. Journal of Statistical Planning and Inference, 227(??):112–128, December 2023. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375823000320.

Kengne:2024:DLW

[KW24] William Kengne and Modou Wade. Deep learning for ψweakly dependent processes. Journal of Statistical Planning and Inference, 232(??):Article 106163, September 2024. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837582400020X.

Kabaila:2020:FSP

[KWW20] Paul Kabaila, A. H. Welsh, and Christeen Wijethunga. Finite sample properties of confidence intervals centered on a model averaged estimator. Journal of Statistical Planning and Inference, 207 (??):10–26, July 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375819300990.

Kojadinovic:2024:RTC

[KY24] Ivan Kojadinovic and Bingqing Yi. Resampling techniques for a class of smooth, possibly data-adaptive empirical copulas. Journal of Statistical Planning and Inference, 231(??):Article 106132, July 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823001015.

Karl:2021:DBL

[KZ21] Andrew T. Karl and Dale L. Zimmerman. A diagnostic for bias in linear mixed model estimators induced by dependence between the random effects and the corresponding model matrix. Journal of Statistical Planning and Inference, 211(??):107– 118, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375820300732.

Kong:2023:RPD

[KZA23] Xiangshun Kong, Wei Zheng, and Mingyao Ai. Representative points for distribution recovering. Journal of Statistical Planning and Inference, 224(??):69–83, May 2023. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837582200091X.

Kong:2023:CDT

[KZZ23] Xiangshun Kong, Xueru Zhang, and Wei Zheng. Circular designs for total effects under interference models. Journal of Statistical Planning and Inference, 227(??):146–161, December 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823000344.

Lane:2023:EOI

[Lan23] Adam Lane. Efficiency of observed information adaptive designs in linear models. Journal of Statistical Planning and Inference, 223(??):63–74, March 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375822000805.

Lando:2022:PIO

[LAO22] Tommaso Lando, Idir Arab, and Paulo Eduardo Oliveira. Properties of increasing odds rate distributions with a statistical application. Journal of Statistical Planning and Inference, 221(??):313– 325, December 2022. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375822000428.

Leach:2022:ISS

[LAYT22] Justin M. Leach, Inmaculada Aban, Nengjun Yi, and The Alzheimerś Disease Neuroimaging Initiative. Incorporating spatial structure into inclusion probabilities for Bayesian variable selection in generalized linear models with the spike-and-slab elastic net. Journal of Statistical Planning and Inference, 217(??): 141–152, March 2022. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837582100077X.

Liao:2022:MPD

[LB22] Yijie Liao and Ronald W. Butler. Marginal posterior distributions for regression parameters in the Cox model using Dirichlet and gamma process priors. Journal of Statistical Planning and Inference, 216(??):95–108, January 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000513.

Liu:2020:GSR

[LC20] Wenyu Liu and D. Stephen Coad. Group-sequential responseadaptive designs for censored survival outcomes. Journal of Statistical Planning and Inference, 205(??):293–305, March 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837581930076X.

Leaver:2024:CLB

[LCKB24] Victoria L. Leaver, Robert G. Clark, Pavel N. Krivitsky, and Carole L. Birrell. A comparison of likelihood-based methods for size-biased sampling. Journal of Statistical Planning and Inference, 230(??):Article 106115, May 2024. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375823000848.

Lee:2022:PKQ

[LCP22] Eun Ryung Lee, Jinwoo Cho, and Seyoung Park. Penalized kernel quantile regression for varying coefficient models. Journal of Statistical Planning and Inference, 217(??):8–23, March 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000707.

Li:2021:GOG

[LCYZ21] Yuanbo Li, Ngai Hang Chan, Chun Yip Yau, and Rongmao Zhang. Group orthogonal greedy algorithm for change-point estimation of multivariate time series. Journal of Statistical Planning and Inference, 212(??):14–33, May 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820301002.

Li:2024:ZVT

[LCZ24] Qi Li, Huaping Chen, and Fukang Zhu. Z-valued time series: Models, properties and comparison. Journal of Statistical Planning and Inference, 230(??):Article 106099, May 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837582300068X.

Liu:2021:SPH

[LFLL21] Guanfu Liu, Yan Fan, Yang Liu, and Yukun Liu. Semiparametric homogeneity test and sample size calculation for a two-sample problem under an inequality constraint. Journal of Statistical Planning and Inference, 214(??):13–24, September 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837582100001X.

Lin:2020:RNA

[LFR20] Zhantao Lin, Nancy Flournoy, and William F. Rosenberger. Random norming aids analysis of non-linear regression models with sequential informative dose selection. Journal of Statistical Planning and Inference, 206(??):29–42, May 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375819300783.

Lopez-Fidalgo:2022:RAL

[LFW22] Jesús López-Fidalgo and Douglas P. Wiens. Robust active learning with binary responses. Journal of Statistical Planning and Inference, 220(??):1–14, September 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375822000040.

Li:2021:BCP

[LG21] Xiuqi Li and Subhashis Ghosal. Bayesian change point detection for functional data. Journal of Statistical Planning and Inference, 213(??):193–205, July 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820301336.

Luo:2023:SBD

[LGM23] Yu Luo, Daniel J. Graham, and Emma J. McCoy. Semiparametric Bayesian doubly robust causal estimation. Journal of Statistical Planning and Inference, 225(??):171–187, July 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822001124.

Lu:2020:NPI

[LJ20] Hezhi Lu and Hua Jin. A new prediction interval for binomial random variable based on inferential models. Journal of Statistical Planning and Inference, 205(??):156–174, March 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819300680.

Liu:2023:FSU

[LL23] Yuyang Liu and Shan Luo. Feature selection in ultrahighdimensional additive models with heterogeneous frequency component functions. Journal of Statistical Planning and Inference, 225(??):243–265, July 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375823000046.

Levy-Leduc:2022:SAE

[LLBR22] Céline Lévy-Leduc, Pascal Bondon, and Valdério Reisen. A spectral approach to estimate the autocovariance function. Journal of Statistical Planning and Inference, 221(??):281–298, December 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837582200043X.

Lloyd:2022:ECL

[Llo22] Chris J. Lloyd. Exact confidence limits compatible with the result of a sequential trial. Journal of Statistical Planning and Inference, 217(??):171–176, March 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000811.

Li:2022:SCC

[LLT22] Wenlong Li, Min-Qian Liu, and Boxin Tang. A systematic construction of compromise designs under baseline parameterization. Journal of Statistical Planning and Inference, 219(??): 33–42, July 2022. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375821001129.

Li:2021:CON

[LLY21] Wenlong Li, Min-Qian Liu, and Jian-Feng Yang. Columnorthogonal nearly strong orthogonal arrays. Journal of Statistical Planning and Inference, 215(??):184–192, December 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000343.

Li:2021:OPQ

[LLZL21] Rui Li, Wenqi Lu, Zhongyi Zhu, and Heng Lian. Optimal prediction of quantile functional linear regression in reproducing kernel Hilbert spaces. Journal of Statistical Planning and Inference, 211(??):162–170, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837582030080X.

Lu:2021:OED

[LMMJ21] Xiao-Nan Lu, Miwako Mishima, Nobuko Miyamoto, and Masakazu Jimbo. Optimal and efficient designs for fMRI experiments via two-level circulant almost orthogonal arrays. Journal of Statistical Planning and Inference, 213(??):33–49, July 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820301191.

Laha:2021:BCD

[LMW21] Nilanjana Laha, Zhen Miao, and Jon A. Wellner. Bi-s∗-concave distributions. Journal of Statistical Planning and Inference, 215(??):127–157, December 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375821000331.

Liu:2023:SRM

[LNH^{+}23] Xiangyu Liu, Jing Ning, Xuming He, Barbara C. Tilley, and Ruosha Li. Semiparametric regression modeling of the global percentile outcome. Journal of Statistical Planning and Inference, 222(??):149–159, January 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375822000623.

Lopes:2020:ESC

[Lop20] Miles E. Lopes. Estimating a sharp convergence bound for randomized ensembles. Journal of Statistical Planning and Inference, 204(??):35–44, January 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375819300369.

Longla:2021:NRC

[LP21] Martial Longla and Magda Peligrad. New robust confidence intervals for the mean under dependence. Journal of Statistical Planning and Inference, 211(??):90–106, March 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375818302416.

Li:2021:SSS

[LPK^{+}21] Ning Li, Xiaoling Peng, Eric Kawaguchi, Marc A. Suchard, and Gang Li. A scalable surrogate L_{0} sparse regression method for generalized linear models with applications to large scale data. Journal of Statistical Planning and Inference, 213(??):262–281, July 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837582030135X.

Lee:2021:VLE

[LPL21] Shen-Ming Lee, Kim-Hung Pho, and Chin-Shang Li. Validation likelihood estimation method for a zero-inflated Bernoulli regression model with missing covariates. Journal of Statistical Planning and Inference, 214(??):105–127, September 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000069.

Liu:2021:NVP

[LTW21] Yin Liu, Guo-Liang Tian, and Mingqiu Wang. A new variant of the parallel regression model with variable selection in surveys with sensitive attribute. Journal of Statistical Planning and Inference, 212(??):69–83, May 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820301129.

Lin:2022:DNI

[LTZF22] Yu-Xuan Lin, Yi-Han Tang, Jia-Hua Zhang, and Kai-Tai Fang. Detecting non-isomorphic orthogonal design. Journal of Statistical Planning and Inference, 221(??):299–312, December 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375822000416.

Lin:2020:EAT

[LW20] Chien-Ju Lin and James M. S. Wason. Efficient analysis of time-to-event endpoints when the event involves a continuous variable crossing a threshold. Journal of Statistical Planning and Inference, 208(??):119–129, September 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820300185.

Lyu:2022:ICS

[LW22] Ziyang Lyu and A. H. Welsh. Increasing cluster size asymptotics for nested error regression models. Journal of Statistical Planning and Inference, 217(??):52–68, March 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000768.

Li:2024:CMD

[LWL^{+}24] Kexuan Li, Fangfang Wang, Ruiqi Liu, Fan Yang, and Zuofeng Shang. Calibrating multi-dimensional complex ODE from noisy data via deep neural networks. Journal of Statistical Planning and Inference, 232(??):Article 106147, September 2024. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375824000041.

Li:2023:TIH

[LX23] Weiming Li and Xiaoge Xiong. A test for the identity of a high-dimensional correlation matrix based on the ℓ_{4}-norm. Journal of Statistical Planning and Inference, 225(??):132–145, July 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822001094.

Lu:2020:SBA

[LY20] Renjie Lu and Philip L. H. Yu. Smooth buffered autoregressive time series models. Journal of Statistical Planning and Inference, 206(??):196–210, May 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837581930093X.

Liu:2021:FTS

[LY21] Zhan Liu and Chun Yip Yau. Fitting time series models for longitudinal surveys with nonignorable missing data. Journal of Statistical Planning and Inference, 214(??):1–12, September 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000021.

Li:2024:COS

[LYCL24] Hui Li, Liuqing Yang, Kashinath Chatterjee, and Min-Qian Liu. Construction of optimal supersaturated designs by the expansive replacement method. Journal of Statistical Planning and Inference, 230(??):Article 106118, May 2024. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375823000873.

Li:2020:REM

[LYH20] Wei Li, Shu Yang, and Peisong Han. Robust estimation for moment condition models with data missing not at random. Journal of Statistical Planning and Inference, 207(??):246–254, July 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818301204.

Liu:2022:USI

[LYH22] Shu Liu, Jinhong You, and Lixia Hu. Unified statistical inference for a nonlinear dynamic functional/longitudinal data model. Journal of Statistical Planning and Inference, 219(??):175–188, July 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821001336.

Li:2021:CLT

[LYZ21] Huiqin Li, Yanqing Yin, and Shurong Zheng. Central limit theorem for linear spectral statistics of general separable sample covariance matrices with applications. Journal of Statistical Planning and Inference, 211(??):80–89, March 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820300720.

Li:2020:ODQ

[LZ20a] Chang Li and Chongqi Zhang. A-optimal designs for quadratic mixture canonical polynomials with spline. Journal of Statistical Planning and Inference, 207(??):1–9, July 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375819300989.

Liu:2020:JEH

[LZ20b] Qingyang Liu and Yuping Zhang. Joint estimation of heterogeneous exponential Markov Random Fields through an approximate likelihood inference. Journal of Statistical Planning and Inference, 209(??):252–266, December 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820300410.

Liu:2022:ECP

[LZ22] Xialu Liu and Ting Zhang. Estimating change-point latent factor models for high-dimensional time series. Journal of Statistical Planning and Inference, 217(??):69–91, March 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000732.

Lu:2024:SMK

[LZLL24] Wenqi Lu, Zhongyi Zhu, Rui Li, and Heng Lian. Sparse multiple kernel learning: Minimax rates with random projection. Journal of Statistical Planning and Inference, 231(??):Article 106142, July 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823001118.

Liang:2024:NAU

[LZW^{+}24] Wanfeng Liang, Yuhao Zhang, Jiyang Wang, Yue Wu, and Xiaoyan Ma. A new approach for ultrahigh dimensional precision matrix estimation. Journal of Statistical Planning and Inference, 232(??):Article 106164, September 2024. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375824000211.

Medina-Aguayo:2022:PWM

[MAC22] Felipe J. Medina-Aguayo and J. Andrés Christen. Penalised t-walk MCMC. Journal of Statistical Planning and Inference, 221(??):230–247, December 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375822000386.

Maitra:2020:CBA

[MB20] Trisha Maitra and Sourabh Bhattacharya. On classical and Bayesian asymptotics in stochastic differential equations with random effects having mixture normal distributions. Journal of Statistical Planning and Inference, 208(??):36–57, September 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300057.

Moutanabbir:2024:NNP

[MB24] Khouzeima Moutanabbir and Mohammed Bouaddi. A new non-parametric estimation of the expected shortfall for dependent financial losses. Journal of Statistical Planning and Inference, 232(??):Article 106151, September 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375824000089.

Mukherjee:2023:CAR

[MCJ23] Ayon Mukherjee, D. Stephen Coad, and Sayantee Jana. Covariateadjusted response-adaptive designs for censored survival responses. Journal of Statistical Planning and Inference, 225(??):219–242, July 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823000010.

Mulgrave:2023:BAN

[MG23] Jami J. Mulgrave and Subhashis Ghosal. Bayesian analysis of nonparanormal graphical models using rank-likelihood. Journal of Statistical Planning and Inference, 222(??):195–208, January 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000593.

Mies:2021:ESD

[Mie21] Fabian Mies. Estimation of state-dependent jump activity and drift for Markovian semimartingales. Journal of Statistical Planning and Inference, 210(??):114–140, January 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820300471.

Mulder:2021:BTC

[MK21] Kees Tim Mulder and Irene Klugkist. Bayesian tests for circular uniformity. Journal of Statistical Planning and Inference, 211(??): 315–325, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837581830329X.

Malcherczyk:2021:SDA

[MLM21] Dennis Malcherczyk, Kevin Leckey, and Christine H. Müller. Ksign depth: From asymptotics to efficient implementation. Journal of Statistical Planning and Inference, 215(??):344–355, December 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000458.

Mondal:2020:SRS

[MM20] Anirban Mondal and Abhijit Mandal. Stratified random sampling for dependent inputs in Monte Carlo simulations from computer experiments. Journal of Statistical Planning and Inference, 205(??):269–282, March 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375819300746.

Meziani:2020:PQP

[MMD20] Aymen Meziani, Tarek Medkour, and Karim Djouani. Penalised quantile periodogram for spectral estimation. Journal of Statistical Planning and Inference, 207(??):86–98, July 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818302489.

McCloud:2021:CDF

[MP21] Nadine McCloud and Christopher F. Parmeter. Calculating degrees of freedom in multivariate local polynomial regression. Journal of Statistical Planning and Inference, 210(??):141–160, January 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300598.

Macci:2023:IST

[MP23] Claudio Macci and Mauro Piccioni. An inverse Sanov theorem for exponential families. Journal of Statistical Planning and Inference, 224(??):54–68, May 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375822000921.

Milhaud:2022:STS

[MPSV22] Xavier Milhaud, Denys Pommeret, Yahia Salhi, and Pierre Vandekerkhove. Semiparametric two-sample admixture components comparison test: the symmetric case. Journal of Statistical Planning and Inference, 216(??):135–150, January 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000616.

McElroy:2021:TAS

[MR21] Tucker McElroy and Anindya Roy. Testing for adequacy of seasonal adjustment in the frequency domain. Journal of Statistical Planning and Inference, 211(??):241–255, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300823.

Marchand:2021:MER

[MRS21] Éric Marchand, Fanny Rancourt, and William E. Strawderman. Minimax estimation of a restricted mean for a oneparameter exponential family. Journal of Statistical Planning and Inference, 212(??):114–125, May 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820301038.

Maruyama:2021:GSA

[MS21a] Yuzo Maruyama and William E. Strawderman. A Gaussian sequence approach for proving minimaxity: a review. Journal of Statistical Planning and Inference, 211(??):256–270, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818303276.

Matsuda:2021:PDE

[MS21b] Takeru Matsuda and William E. Strawderman. Predictive density estimation under the Wasserstein loss. Journal of Statistical Planning and Inference, 210(??):53–63, January 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837582030063X.

Maruyama:2023:RBA

[MS23] Yuzo Maruyama and William E. Strawderman. A review of Brown 1971 (in)admissibility results under scale mixtures of Gaussian priors. Journal of Statistical Planning and Inference, 222(??):78– 93, January 2023. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375822000520.

Mroz:2023:DFM

[MSFT23] Thomas Mroz, Juan Fernández Sánchez, Sebastian Fuchs, and Wolfgang Trutschnig. On distributions with fixed marginals maximizing the joint or the prior default probability, estimation, and related results. Journal of Statistical Planning and Inference, 223(??):33–52, March 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375822000684.

Mori:2021:ETN

[MSR21] Tamás F. Móri, Gábor J. Székely, and Maria L. Rizzo. On energy tests of normality. Journal of Statistical Planning and Inference, 213(??):1–15, July 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375820301154.

Matsuoka:2023:NES

[MT23] Danilo H. Matsuoka and Hudson S. Torrent. Nonparametric estimation of a smooth trend in the presence of a periodic sequence. Journal of Statistical Planning and Inference, 225(??): 202–218, July 2023. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375823000022.

Masuda:2021:EDC

[MU21] Hiroki Masuda and Yuma Uehara. Estimating diffusion with compound Poisson jumps based on self-normalized residuals. Journal of Statistical Planning and Inference, 215(??):158–183, December 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000240.

Ma:2024:DDR

[MWL24] Xiaoyan Ma, Lili Wei, and Wanfeng Liang. D4R: Doubly robust reduced rank regression in high dimension. Journal of Statistical Planning and Inference, 232(??):Article 106162, September 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375824000193.

Narukawa:2021:ETL

[Nar21] Masaki Narukawa. Efficient tapered local Whittle estimation of multivariate fractional processes. Journal of Statistical Planning and Inference, 215(??):234–256, December 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000380.

Neumann:2021:EPT

[NBD21] André Neumann, Taras Bodnar, and Thorsten Dickhaus. Estimating the proportion of true null hypotheses under dependency: a marginal bootstrap approach. Journal of Statistical Planning and Inference, 210(??):76–86, January 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820300495.

Ninomiya:2021:MMP

[NKST21] Yoshiyuki Ninomiya, Satoshi Kuriki, Toshihiko Shiroishi, and Toyoyuki Takada. A modification of MaxT procedure using spurious correlations. Journal of Statistical Planning and Inference, 214(??):128–138, September 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000173.

Ng:2020:LTC

[NLL20] Chi Tim Ng, Woojoo Lee, and Youngjo Lee. Logical and test consistency in pairwise multiple comparisons. Journal of Statistical Planning and Inference, 206(??):145–162, May 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818301290.

Nadarajah:2021:OBC

[NMP21] K. Nadarajah, Gael M. Martin, and D. S. Poskitt. Optimal bias correction of the log-periodogram estimator of the fractional parameter: a jackknife approach. Journal of Statistical Planning and Inference, 211(??):41–79, March 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820300483.

Novikov:2020:OMS

[NRP20] Andrey Novikov and Pedro Reyes-Pérez. Optimal multistage sequential hypothesis testing. Journal of Statistical Planning and Inference, 205(??):219–230, March 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375819300710.

Nandi:2024:FRC

[NS24] Shinjini Nandi and Sanat K. Sarkar. Further results on controlling the false discovery rate under some complex grouping structure of hypotheses. Journal of Statistical Planning and Inference, 229(??):Article 106094, March 2024. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375823000551.

Nandi:2021:AOT

[NSC21] Shinjini Nandi, Sanat K. Sarkar, and Xiongzhi Chen. Adapting to one- and two-way classified structures of hypotheses while controlling the false discovery rate. Journal of Statistical Planning and Inference, 215(??):95–108, December 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000227.

Noonan:2022:RQR

[NZ22] Jack Noonan and Anatoly Zhigljavsky. Random and quasirandom designs in group testing. Journal of Statistical Planning and Inference, 221(??):29–54, December 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375822000180.

Ouimet:2021:PLL

[Oui21] Frédéric Ouimet. A precise local limit theorem for the multinomial distribution and some applications. Journal of Statistical Planning and Inference, 215(??):218–233, December 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000392.

Overstall:2022:PFI

[Ove22] Antony M. Overstall. Properties of Fisher information gain for Bayesian design of experiments. Journal of Statistical Planning and Inference, 218(??):138–146, May 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821001087.

Ohishi:2020:FAO

[OYF20] Mineaki Ohishi, Hirokazu Yanagihara, and Yasunori Fujikoshi. A fast algorithm for optimizing ridge parameters in a generalized ridge regression by minimizing a model selection criterion. Journal of Statistical Planning and Inference, 204(??):187–205, January 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819300503.

Patriota:2022:MFM

[PdOA22] Alexandre Galvão Patriota and Jônatas de Oliveira Alves. A monotone frequentist measure of evidence for testing variance components in linear mixed models. Journal of Statistical Planning and Inference, 219(??):43–62, July 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821001105.

Penda:2024:KET

[Pen24] S. Valère Bitseki Penda. Kernel estimation of the transition density in bifurcating Markov chains. Journal of Statistical Planning and Inference, 231(??):Article 106138, July 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823001076.

Pinheiro:2021:HOA

[PFM21] Eliane C. Pinheiro, Silvia L. P. Ferrari, and Francisco M. C. Medeiros. Higher-order approximate confidence intervals. Journal of Statistical Planning and Inference, 213(??):206–221, July 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820301348.

Pilz:2020:CEN

[Pil20] Maximilian Pilz. Consistency of the Elastic Net under a finite second moment assumption on the noise. Journal of Statistical Planning and Inference, 204(??):72–79, January 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375818302210.

Peng:2020:SSD

[PL20] Jiayu Peng and Dennis K. J. Lin. Small screening design when the overall variance is unknown. Journal of Statistical Planning and Inference, 205(??):1–9, March 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375819300515.

Park:2020:SIM

[PPTO20] Hyung Park, Eva Petkova, Thaddeus Tarpey, and R. Todd Ogden. A single-index model with multiple-links. Journal of Statistical Planning and Inference, 205(??):115–128, March 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818301678.

Proia:2020:MDC

[Pro20] Frédéric Proïa. Moderate deviations in a class of stable but nearly unstable processes. Journal of Statistical Planning and Inference, 208(??):66–81, September 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820300070.

Prus:2022:ETM

[Pru22] Maryna Prus. Equivalence theorems for multiple-design problems with application in mixed models. Journal of Statistical Planning and Inference, 217(??):153–164, March 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000793.

Pronzato:2021:RIP

[PS21] Luc Pronzato and Guillaume Sagnol. Removing inessential points in c-and A-optimal design. Journal of Statistical Planning and Inference, 213(??):233–252, July 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820301324.

Patilea:2020:TLF

[PSS20] Valentin Patilea and César Sánchez-Sellero. Testing for lack-of-fit in functional regression models against general alternatives. Journal of Statistical Planning and Inference, 209(??):229–251, December 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375820300409.

Pronzato:2021:SOS

[PW21] Luc Pronzato and HaiYing Wang. Sequential online subsampling for thinning experimental designs. Journal of Statistical Planning and Inference, 212(??):169–193, May 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820300999.

Pournaderi:2024:VSK

[PX24] Mehrdad Pournaderi and Yu Xiang. Variable selection with the knockoffs: Composite null hypotheses. Journal of Statistical Planning and Inference, 231(??):Article 106119, July 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823000885.

Qiao:2024:OSC

[QLXL24] Nan Qiao, Wangcheng Li, Feng Xiao, and Cunjie Lin. Optimal subsampling for the Cox proportional hazards model with massive survival data. Journal of Statistical Planning and Inference, 231(??):Article 106136, July 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375823001052.

Qi:2023:DCM

[QLZ23] Liangwei Qi, Ze Liu, and Yongdao Zhou. Deterministic construction methods for uniform designs. Journal of Statistical Planning and Inference, 226(??):30–38, September 2023. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375823000137.

Qiu:2023:GAF

[QW23] Zhiping Qiu and Dixin Wu. Generalized accelerated failure time model with censored data from case-cohort studies. Journal of Statistical Planning and Inference, 224(??):42–53, May 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000891.

Rigon:2021:TBD

[RD21] Tommaso Rigon and Daniele Durante. Tractable Bayesian density regression via logit stick-breaking priors. Journal of Statistical Planning and Inference, 211(??):131–142, March 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375818300697.

Rezende:2022:NFI

[Rez22] Thiago Rezende. A new filtering inference procedure for a GED state-space volatility model. Journal of Statistical Planning and Inference, 221(??):55–68, December 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375822000210.

Rademacher:2024:SIW

[RKvS24] Daniel Rademacher, Johannes Krebs, and Rainer von Sachs. Statistical inference for wavelet curve estimators of symmetric positive definite matrices. Journal of Statistical Planning and Inference, 231(??):Article 106140, July 2024. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837582300109X.

Rosa:2020:OED

[Ros20] Samuel Rosa. Optimal experimental designs for treatment contrasts in heteroscedastic models with covariates. Journal of Statistical Planning and Inference, 209(??):280–291, December 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300434.

Ruli:2020:RAB

[RSV20] Erlis Ruli, Nicola Sartori, and Laura Ventura. Robust approximate Bayesian inference. Journal of Statistical Planning and Inference, 205(??):10–22, March 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837581830106X.

Rempala:2023:PLT

[RW23] Grzegorz Rempała and Jacek Wesołowski. Poisson limit theorems for the Cressie–Read statistics. Journal of Statistical Planning and Inference, 223(??):15–32, March 2023. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375822000672.

Reiss:2020:TPS

[RX20] Philip T. Reiss and Meng Xu. Tensor product splines and functional principal components. Journal of Statistical Planning and Inference, 208(??):1–12, September 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375819301132.

Ryabko:2022:AMP

[Rya22] Boris Ryabko. Asymptotically most powerful tests for random number generators. Journal of Statistical Planning and Inference, 217(??):1–7, March 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375821000744.

Stocker:2020:GCA

[SA20] Russell Stocker and Akim Adekpedjou. A general class of additive semiparametric models for recurrent event data. Journal of Statistical Planning and Inference, 205(??):231–244, March 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819300722.

Sabbaghi:2021:IFG

[Sab21] Arman Sabbaghi. An integrative framework for geometric and hidden projections in three-level fractional factorial designs. Journal of Statistical Planning and Inference, 215(??):257–267, December 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000367.

Saegusa:2021:NID

[Sae21] Takumi Saegusa. Nonparametric inference for distribution functions with stratified samples. Journal of Statistical Planning and Inference, 215(??):356–367, December 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837582100046X.

Saegusa:2022:SIM

[Sae22] Takumi Saegusa. Semiparametric inference for merged data from multiple data sources. Journal of Statistical Planning and Inference, 216(??):1–14, January 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000471.

Salamanca:2021:UCD

[Sal21] Juan Jesús Salamanca. A universal, canonical dispersive ordering in metric spaces. Journal of Statistical Planning and Inference, 212 (??):1–13, May 2021. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375820301117.

Sattler:2022:THA

[SBP22] Paavo Sattler, Arne C. Bathke, and Markus Pauly. Testing hypotheses about covariance matrices in general MANOVA designs. Journal of Statistical Planning and Inference, 219(??): 134–146, July 2022. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375821001269.

Sun:2024:RAL

[SCCF24] Zhuowei Sun, Hongyuan Cao, Li Chen, and Jason P. Fine. Regression analysis of longitudinal data with mixed synchronous and asynchronous longitudinal covariates. Journal of Statistical Planning and Inference, 231(??):Article 106135, July 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823001040.

Soale:2021:EAI

[SD21] Abdul-Nasah Soale and Yuexiao Dong. On expectile-assisted inverse regression estimation for sufficient dimension reduction. Journal of Statistical Planning and Inference, 213(??):80–92, July 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837582030118X.

Shi:2021:CAF

[SDSZ21] Gongming Shi, Jiang Du, Zhihua Sun, and Zhongzhan Zhang. Checking the adequacy of functional linear quantile regression model. Journal of Statistical Planning and Inference, 210(??):64– 75, January 2021. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375820300616.

Singh:2021:OFD

[SKS21] Rakhi Singh, Joachim Kunert, and John Stufken. On optimal fMRI designs for correlated errors. Journal of Statistical Planning and Inference, 212(??):84–96, May 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820301014.

Sun:2020:NCM

[SLJ20] Hanmei Sun, Yihui Luan, and Jiming Jiang. A new classified mixed model predictor. Journal of Statistical Planning and Inference, 207 (??):45–54, July 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375819301028.

Sun:2024:SAF

[SLW24] Yifan Sun, Ziyi Liu, and Wu Wang. Subgroup analysis for the functional linear model. Journal of Statistical Planning and Inference, 231(??):Article 106120, July 2024. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375823000897.

Szabo:2020:SSM

[SLX20] Zsolt Szabo, Xiaoyu Liu, and Liming Xiang. Semiparametric sieve maximum likelihood estimation for accelerated hazards model with interval-censored data. Journal of Statistical Planning and Inference, 205(??):175–192, March 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375818302477.

Singh:2023:SBO

[SM23] Rahul Singh and Neeraj Misra. U-statistic based on overlapping sample spacings. Journal of Statistical Planning and Inference, 224 (??):98–108, May 2023. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375822000945.

Salles:2020:SEE

[SMB20] Gabriel Salles, Sophie Mercier, and Laurent Bordes. Semiparametric estimate of the efficiency of imperfect maintenance actions for a gamma deteriorating system. Journal of Statistical Planning and Inference, 206(??):278–297, May 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375819300977.

She:2022:WPE

[SML22] Rui She, Zichuan Mi, and Shiqing Ling. Whittle parameter estimation for vector ARMA models with heavy-tailed noises. Journal of Statistical Planning and Inference, 219(??):216–230, July 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821001257.

Schmidt:2020:ODP

[SS20] Marius Schmidt and Rainer Schwabe. Optimal designs for Poisson count data with Gamma block effects. Journal of Statistical Planning and Inference, 204(??):128–140, January 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375818302039.

Sofro:2020:RAM

[SSC20] A’yunin Sofro, Jian Qing Shi, and Chunzheng Cao. Regression analysis for multivariate process data of counts using convolved Gaussian processes. Journal of Statistical Planning and Inference, 206(??):57–74, May 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375819300801.

Scholtus:2022:CLE

[SSdW22] Sander Scholtus, Natalie Shlomo, and Ton de Waal. Correcting for linkage errors in contingency tables — a cautionary tale. Journal of Statistical Planning and Inference, 218(??):122–137, May 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821001063.

Shevchenko:2020:GVH

[SST20] Radomyra Shevchenko, Meryem Slaoui, and C. A. Tudor. Generalized k-variations and Hurst parameter estimation for the fractional wave equation via Malliavin calculus. Journal of Statistical Planning and Inference, 207(??):155–180, July 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375819301193.

Song:2021:NMH

[ST21] Xiaojun Song and Abderrahim Taamouti. A nonparametric measure of heteroskedasticity. Journal of Statistical Planning and Inference, 212(??):45–68, May 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375820301105.

Stein:2022:NSD

[Ste22] Michael L. Stein. A note on the smoothness of densities. Journal of Statistical Planning and Inference, 217(??):165–170, March 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000884.

Su:2022:ASS

[Su22] Yingcai Su. On the admissibility of spherical spatial covariance functions in higher dimensions. Journal of Statistical Planning and Inference, 217(??):92–105, March 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000756.

Sun:2022:AGL

[SW22] Yifan Sun and Qihua Wang. An adaptive group LASSO approach for domain selection in functional generalized linear models. Journal of Statistical Planning and Inference, 219(??):13–32, July 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821001117.

Sheng:2024:NFO

[SWS24] Danshu Sheng, Dehui Wang, and Liuquan Sun. A new first-order mixture integer-valued threshold autoregressive process based on binomial thinning and negative binomial thinning. Journal of Statistical Planning and Inference, 231(??):Article 106143, July 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837582300112X.

Satter:2021:JEL

[SZ21] Faysal Satter and Yichuan Zhao. Jackknife empirical likelihood for the mean difference of two zero-inflated skewed populations. Journal of Statistical Planning and Inference, 211(??):414–422, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300914.

Su:2022:TAS

[SZ22] Bing Su and Fukang Zhu. Temporal aggregation and systematic sampling for INGARCH processes. Journal of Statistical Planning and Inference, 219(??):120–133, July 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821001270.

Sun:2023:GML

[SZ23] Tao Sun and Shengli Zhao. General minimum lower-order confounding split-plot designs with important whole-plot factors. Journal of Statistical Planning and Inference, 224(??):1–8, May 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000829.

Tan:2022:AOP

[Tan22] Zhiqiang Tan. Analysis of odds, probability, and hazard ratios: From 2 by 2 tables to two-sample survival data. Journal of Statistical Planning and Inference, 221(??):248–265, December 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000404.

Tyssedal:2022:PPP

[TH22] John Tyssedal and Yngvild Hole Hamre. Preserving projection properties when regular two-level designs are blocked. Journal of Statistical Planning and Inference, 221(??):266–280, December 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000398.

Torabi:2020:EMS

[TJ20] Mahmoud Torabi and Jiming Jiang. Estimation of mean squared prediction error of empirically spatial predictor of small area means under a linear mixed model. Journal of Statistical Planning and Inference, 208(??):82–93, September 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820300161.

Tian:2024:ICS

[TL24] Zibo Tian and Fan Li. Information content of stepped wedge designs under the working independence assumption. Journal of Statistical Planning and Inference, 229(??):Article 106097, March 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823000666.

Turner:2024:GVE

[TLG24] Rosanne J. Turner, Alexander Ly, and Peter D. Grünwald. Generic e-variables for exact sequential k-sample tests that allow for optional stopping. Journal of Statistical Planning and Inference, 230(??):Article 106116, May 2024. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837582300085X.

Tung:2022:AAD

[TLT22] Hung-Ping Tung, I-Chen Lee, and Sheng-Tsaing Tseng. Analytical approach for designing accelerated degradation tests under an exponential dispersion model. Journal of Statistical Planning and Inference, 218(??):73–84, May 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837582100104X.

Tran:2023:LPJ

[TN23] Ngoc Khue Tran and Hoang-Long Ngo. LAMN property for jump diffusion processes with discrete observations on a fixed time interval. Journal of Statistical Planning and Inference, 225(??): 1–28, July 2023. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375822000970.

Tang:2022:TIP

[TPZ22] Yinfen Tang, Ye Peng, and Zhiyuan Zhang. Trading information, price discreteness, and volatility estimation. Journal of Statistical Planning and Inference, 220(??):49–70, September 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000118.

Telschow:2022:SCB

[TS22] Fabian J. E. Telschow and Armin Schwartzman. Simultaneous confidence bands for functional data using the Gaussian kinematic formula. Journal of Statistical Planning and Inference, 216(??): 70–94, January 2022. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375821000598.

Tsai:2022:GOO

[Tsa22] Shin-Fu Tsai. Generating optimal order-of-addition designs with flexible run sizes. Journal of Statistical Planning and Inference, 218(??):147–163, May 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375821001099.

Trong:2021:DEM

[TTML21] Dang Duc Trong, Nguyen Hoang Thanh, Nguyen Dang Minh, and Nguyen Nhu Lan. Density estimation of a mixture distribution with unknown point-mass and normal error. Journal of Statistical Planning and Inference, 215(??):268–288, December 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000410.

Tong:2020:MQR

[TW20] Hongzhi Tong and Qiang Wu. Moving quantile regression. Journal of Statistical Planning and Inference, 205(??):46–63, March 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581930062X.

Uto:2022:CRG

[UB22] N. P. Uto and R. A. Bailey. Constructions for regulargraph semi-Latin rectangles with block size two. Journal of Statistical Planning and Inference, 221(??):81–89, December 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000192.

Vidal:2021:BAM

[VdC21] Ignacio Vidal and Mário de Castro. A Bayesian analysis of the matching problem. Journal of Statistical Planning and Inference, 212(??):194–200, May 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820301063.

Walter:2023:ACN

[Wal23] Benjamin Walter. Analysis of convolutional neural network image classifiers in a hierarchical max-pooling model with additional local pooling. Journal of Statistical Planning and Inference, 224(??): 109–126, May 2023. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375822000957.

Willems:2024:UVI

[WC24] Pieter Willems and Gerda Claeskens. Uniformly valid inference for partially linear high-dimensional single-index models. Journal of Statistical Planning and Inference, 229(??):Article 106091, March 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823000526.

Wang:2022:SSL

[WCL22] Xiaodi Wang, Xueping Chen, and Dennis K. J. Lin. Sliced symmetrical Latin hypercube designs. Journal of Statistical Planning and Inference, 218(??):59–72, May 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000963.

Wang:2023:CDA

[WD23] Yunlong Wang and Lilun Du. Change-detection-assisted multiple testing for spatiotemporal data. Journal of Statistical Planning and Inference, 227(??):57–74, December 2023. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375823000290.

Weng:2020:ECB

[WF20] Haolei Weng and Yang Feng. On the estimation of correlation in a binary sequence model. Journal of Statistical Planning and Inference, 207(??):123–137, July 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375819301168.

Wei:2020:CPS

[WG20] Ran Wei and Subhashis Ghosal. Contraction properties of shrinkage priors in logistic regression. Journal of Statistical Planning and Inference, 207(??):215–229, July 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375819301235.

Weller:2020:NSD

[WH20] Zachary D. Weller and Jennifer A. Hoeting. A nonparametric spectral domain test of spatial isotropy. Journal of Statistical Planning and Inference, 204(??):177–186, January 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819300497.

Wechsung:2023:SIP

[WK23] Maximilian Wechsung and Frank Konietschke. Simultaneous inference for partial areas under receiver operating curves — with a view towards efficiency. Journal of Statistical Planning and Inference, 226(??):71–85, September 2023. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375823000149.

Wu:2020:PTH

[WL20] Tung-Lung Wu and Ping Li. Projected tests for high-dimensional covariance matrices. Journal of Statistical Planning and Inference, 207(??):73–85, July 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375818302817.

Wang:2023:ELS

[WL23] Bao-Hua Wang and Han-Ying Liang. Empirical likelihood in single-index quantile regression with high dimensional and missing observations. Journal of Statistical Planning and Inference, 226(??):1–19, September 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375823000058.

Wang:2024:UMP

[WL24] Yining Wang and Gang Li. Uniformly more powerful tests for a subset of the components of a Normal Mean Vector. Journal of Statistical Planning and Inference, 231(??):Article 106141, July 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823001106.

Wang:2024:LAS

[WLFL24] Yue Wang, Hongmei Lin, Zengyan Fan, and Heng Lian. Locally adaptive sparse additive quantile regression model with TV penalty. Journal of Statistical Planning and Inference, 232(??):Article 106144, September 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375824000016.

Wang:2020:MFP

[WM20] Zhe Wang and Ryan Martin. Model-free posterior inference on the area under the receiver operating characteristic curve. Journal of Statistical Planning and Inference, 209(??):174–186, December 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300379.

Wang:2022:SSO

[WM22] Chunyan Wang and Robert W. Mee. Saturated and supersaturated order-of-addition designs. Journal of Statistical Planning and Inference, 219(??):204–215, July 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821001300.

Wang:2024:FCE

[WM24] Chunyan Wang and Robert W. Mee. Fast construction of efficient two-level parallel flats designs. Journal of Statistical Planning and Inference, 229(??):Article 106093, March 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837582300054X.

Wang:2020:FSU

[WMZ20] Lei Wang, Xuejun Ma, and Jingxiao Zhang. Feature screening for ultrahigh-dimensional additive logistic models. Journal of Statistical Planning and Inference, 205(??):306–317, March 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819300771.

Wang:2021:REN

[WST21] WenWu Wang, Wei Shen, and Tiejun Tong. Robust estimation of nonparametric function via addition sequence. Journal of Statistical Planning and Inference, 211(??):423–438, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300896.

Wang:2024:MTG

[WT24] Pengfei Wang and Zhaofeng Tian. Multiple testing in genome-wide association studies via hierarchical hidden Markov models. Journal of Statistical Planning and Inference, 232(??):Article 106161, September 2024. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375824000181.

Wang:2020:ODA

[WVJ^{+}20] Jiali Wang, Arunas P. Verbyla, Bomin Jiang, Alexander B. Zwart, Cheng Soon Ong, Xavier R. R. Sirault, and Klara L. Verbyla. Optimal design for adaptive smoothing splines. Journal of Statistical Planning and Inference, 206(??):263–277, May 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819300965.

Wang:2020:QRP

[WWZZ20] Weiwei Wang, Xianyi Wu, Xiaobing Zhao, and Xian Zhou. Quantile regression for panel count data based on quadratic inference functions. Journal of Statistical Planning and Inference, 207(??): 230–245, July 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375819301247.

Wu:2021:SEA

[WXT^{+}21] Peng Wu, Xinyi Xu, Xingwei Tong, Qing Jiang, and Bo Lu. Semiparametric estimation for average causal effects using propensity score-based spline. Journal of Statistical Planning and Inference, 212(??):153–168, May 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820301099.

Wu:2020:PEV

[WY20] Wenbo Wu and Xiangrong Yin. Pseudo estimation and variable selection in regression. Journal of Statistical Planning and Inference, 208(??):25–35, September 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820300045.

Wang:2022:DOA

[WY22] Jing Wang and Rong-Xian Yue. D-optimal augmented designs and the existence of tight orthogonal arrays with high strength. Journal of Statistical Planning and Inference, 218(??):1–8, May 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000926.

Wang:2024:ISU

[WY24] Lichun Wang and Yang Yang. Inference for seemingly unrelated linear mixed models. Journal of Statistical Planning and Inference, 228(??):80–94, January 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375823000496.

Wang:2021:CSF

[WYL21] Chunyan Wang, Jinyu Yang, and Min-Qian Liu. Construction of space-filling orthogonal designs. Journal of Statistical Planning and Inference, 213(??):130–141, July 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820301245.

Wang:2021:SEP

[WZC^{+}21] Yijun Wang, Jiajia Zhang, Chao Cai, Wenbin Lu, and Yincai Tang. Semiparametric estimation for proportional hazards mixture cure model allowing non-curable competing risk. Journal of Statistical Planning and Inference, 211(??):171–189, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300793.

Wang:2020:DRF

[WZL20] Guochang Wang, Fode Zhang, and Heng Lian. Directional regression for functional data. Journal of Statistical Planning and Inference, 204(??):1–17, January 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375818301228.

Wang:2022:BLT

[WZL^{+}22] Xiaogang Wang, Xiaoying Zhou, Bing Li, Feipeng Zhang, and Xian Zhou. A bent line tobit regression model with application to household financial assets. Journal of Statistical Planning and Inference, 221(??):69–80, December 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375822000209.

Wang:2024:ARH

[WZSQ24] Yinzhi Wang, Yingqiu Zhu, Qiang Sun, and Lei Qin. Adaptively robust high-dimensional matrix factor analysis under Huber loss function. Journal of Statistical Planning and Inference, 231(??):Article 106137, July 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375823001064.

Xu:2020:MDD

[XC20] Kai Xu and Fangxue Chen. Martingale-difference-divergence-based tests for goodness-of-fit in quantile models. Journal of Statistical Planning and Inference, 207(??):138–154, July 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837581930117X.

Xu:2020:EIP

[XDZL20] Wenchao Xu, Hui Ding, Riquan Zhang, and Hua Liang. Estimation and inference in partially functional linear regression with multiple functional covariates. Journal of Statistical Planning and Inference, 209(??):44–61, December 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820300227.

Xu:2020:RMH

[XLZ20] Xinyi Xu, Xiangjie Li, and Jingxiao Zhang. Regularization methods for high-dimensional sparse control function models. Journal of Statistical Planning and Inference, 206(??):111–126, May 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819300904.

Xu:2022:TSF

[XLZL22] Wenchao Xu, Hongmei Lin, Riquan Zhang, and Hua Liang. Twosample functional linear models with functional responses. Journal of Statistical Planning and Inference, 218(??):85–105, May 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821001038.

Xu:2021:IVE

[XMO21] Xiaoming Xu, Mary C. Meyer, and Jean D. Opsomer. Improved variance estimation for inequality-constrained domain mean estimators using survey data. Journal of Statistical Planning and Inference, 215(??):47–71, December 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000203.

Xu:2021:BSA

[XMX21] Jin Xu, Rongji Mu, and Cui Xiong. A Bayesian stochastic approximation method. Journal of Statistical Planning and Inference, 211(??):391–401, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375820300884.

Xu:2021:RDG

[XS21] Xiaojian Xu and Sanjoy K. Sinha. Robust designs for generalized linear mixed models with possible model misspecification. Journal of Statistical Planning and Inference, 210(??):20– 41, January 2021. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375820300446.

Xu:2021:MTT

[XZ21] Kai Xu and Yeqing Zhou. Maximum-type tests for highdimensional regression coefficients using Wilcoxon scores. Journal of Statistical Planning and Inference, 211(??):221–240, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300811.

Yang:2020:RDR

[Yan20] Hojin Yang. Random distributional response model based on spline method. Journal of Statistical Planning and Inference, 207(??): 27–44, July 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375819301016.

Choi:2023:OWP

[yCjL23] Jin young Choi and Myoung jae Lee. Overlap weight and propensity score residual for heterogeneous effects: a review with extensions. Journal of Statistical Planning and Inference, 222(??): 22–37, January 2023. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375822000337.

Yu:2021:EDE

[YCL21] Lili Yu, Ding-Geng Chen, and Jun Liu. Efficient and direct estimation of the variance-covariance matrix in EM algorithm with interpolation method. Journal of Statistical Planning and Inference, 211(??):119–130, March 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820300744.

Yuan:2022:LRT

[YFS22] Mingao Yuan, Yang Feng, and Zuofeng Shang. A likelihoodratio type test for stochastic block models with bounded degrees. Journal of Statistical Planning and Inference, 219(??): 98–119, July 2022. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375821001294.

Yamada:2023:HDA

[YH23] Takayuki Yamada and Tetsuto Himeno. High-dimensional asymptotic expansion of the null distribution for L2 norm based MANOVA testing statistic under general distribution. Journal of Statistical Planning and Inference, 224(??):9–26, May 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000817.

Ye:2022:NPP

[YHDB22] Keying Ye, Zifei Han, Yuyan Duan, and Tianyu Bai. Normalized power prior Bayesian analysis. Journal of Statistical Planning and Inference, 216(??):29–50, January 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000501.

Yin:2020:MFT

[Yin20] Yanqing Yin. Model-free tests for series correlation in multivariate linear regression. Journal of Statistical Planning and Inference, 206(??):179–195, May 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375819300928.

Yang:2024:MCC

[YJL24] Lianqiang Yang, Ying Jing, and Teng Li. Maximum correntropy criterion regression models with tending-to-zero scale parameters. Journal of Statistical Planning and Inference, 231(??):Article 106134, July 2024. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375823001039.

Yuasa:2023:GBE

[YK23] Ryota Yuasa and Tatsuya Kubokawa. Generalized Bayes estimators with closed forms for the normal mean and covariance matrices. Journal of Statistical Planning and Inference, 222(??):182– 194, January 2023. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837582200060X.

Yanagimoto:2024:PNP

[YM24] Takemi Yanagimoto and Yoichi Miyata. A pair of novel priors for improving and extending the conditional MLE. Journal of Statistical Planning and Inference, 231(??):Article 106117, July 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823000861.

Yoon:2020:EIC

[YW20] Hwan-Jin Yoon and Alan H. Welsh. On the effect of ignoring correlation in the covariates when fitting linear mixed models. Journal of Statistical Planning and Inference, 204(??):18– 34, January 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375819300333.

Yu:2020:BFO

[YWYL20] Meiju Yu, Dehui Wang, Kai Yang, and Yan Liu. Bivariate first-order random coefficient integer-valued autoregressive processes. Journal of Statistical Planning and Inference, 204(??):153– 176, January 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375819300485.

Yang:2023:IPM

[YX23] Yuehan Yang and Siwei Xia. An integrated precision matrix estimation for multivariate regression problems. Journal of Statistical Planning and Inference, 222(??):261–272, January 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000659.

Yang:2023:MCP

[YYW23] Xin Yang, Qijing Yan, and Mixia Wu. Model checking for parametric single-index models with massive datasets. Journal of Statistical Planning and Inference, 227(??):129–145, December 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375823000332.

Yang:2021:OSD

[YYYW21] Guijun Yang, Jingbo Yang, Xue Yang, and Weizhen Wang. On optimal subset designs for phase II clinical trials with both total response and disease control. Journal of Statistical Planning and Inference, 215(??):85–94, December 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000239.

Yao:2023:MCI

[YZW23] Yaqiong Yao, Jiahui Zou, and HaiYing Wang. Model constraints independent optimal subsampling probabilities for softmax regression. Journal of Statistical Planning and Inference, 225(??): 188–201, July 2023. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375822001112.

Yi:2022:ODM

[YZZ22] Si-Yu Yi, Yong-Dao Zhou, and Wei Zheng. Optimal designs for mean-covariance models with missing observations. Journal of Statistical Planning and Inference, 219(??):85–97, July 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821001282.

Zhang:2020:PEM

[ZC20a] Jia Zhang and Xin Chen. Principal envelope model. Journal of Statistical Planning and Inference, 206(??):249–262, May 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819300953.

Zhou:2020:ORC

[ZC20b] Yang Zhou and Di-Rong Chen. The optimal rate of canonical correlation analysis for stochastic processes. Journal of Statistical Planning and Inference, 207(??):276–287, July 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837582030001X.

Zhong:2021:CAI

[ZCQC21] Xiaobo Zhong, Ying Kuen Cheung, Min Qian, and Bin Cheng. Comparing adaptive interventions under a general sequential multiple assignment randomized trial design via multiple comparisons with the best. Journal of Statistical Planning and Inference, 211(??):143–153, March 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820300781.

Zeng:2021:GPL

[ZDW^{+}21] JingJing Zeng, Yuze Duan, Desheng Wang, Bin Zou, Yue Yin, and Jie Xu. Generalization performance of Lagrangian support vector machine based on Markov sampling. Journal of Statistical Planning and Inference, 214(??):89–104, September 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300987.

Zou:2020:QRV

[ZFZ20] Yuye Zou, Guoliang Fan, and Riquan Zhang. Quantile regression and variable selection for partially linear single-index models with missing censoring indicators. Journal of Statistical Planning and Inference, 204(??):80–95, January 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375819300382.

Zarepour:2023:QBC

[ZH23] Mahmoud Zarepour and Reza Habibi. A quasi-Bayesian change point detection with exchangeable weights. Journal of Statistical Planning and Inference, 222(??):226–240, January 2023. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375822000441.

Zhang:2021:TSO

[Zha21] Shibin Zhang. A test for second-order stationarity of a time series based on the maximum of Anderson–Darling statistics. Journal of Statistical Planning and Inference, 215(??):109–126, December 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375821000264.

Zhang:2020:MPT

[ZHB20] Qiuyan Zhang, Jiang Hu, and Zhidong Bai. Modified Pillai’s trace statistics for two high-dimensional sample covariance matrices. Journal of Statistical Planning and Inference, 207(??): 255–275, July 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837581830301X.

Zhang:2020:DRK

[ZHLT20] Chenguang Zhang, Hua He, Jian Li, and Wan Tang. Doubly robust kernel density estimation when group membership is missing at random. Journal of Statistical Planning and Inference, 206(??): 163–178, May 2020. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375818302532.

Zhao:2022:FUS

[ZLHZ22] Fanrong Zhao, Nan Lin, Wenjuan Hu, and Baoxue Zhang. A faster u-statistic for testing independence in the functional linear models. Journal of Statistical Planning and Inference, 217(??):188– 203, March 2022. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375821000896.

Zhang:2024:ELB

[ZLW^{+}24] Jing Zhang, Bo Li, Yu Wang, Xinyi Wei, and Xiaohui Liu. An empirical likelihood-based unified test for the integervalued AR(1) models. Journal of Statistical Planning and Inference, 232(??):Article 106149, September 2024. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375824000065.

Zhang:2020:OUC

[ZLZ20] Xue-Ru Zhang, Min-Qian Liu, and Yong-Dao Zhou. Orthogonal uniform composite designs. Journal of Statistical Planning and Inference, 206(??):100–110, May 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375819300898.

Zhou:2022:PTT

[ZMZ22] Mo Zhou, Yaolan Ma, and Rongmao Zhang. Portmanteau-type test for unit root with heavy-tailed noise. Journal of Statistical Planning and Inference, 218(??):25–42, May 2022. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000951.

Zhang:2023:GFT

[ZSS23] Tong Zhang, Zhihua Sun, and Liuquan Sun. Goodness-of-fit test for partial functional linear model with errors in scalar covariates. Journal of Statistical Planning and Inference, 227(??):91– 111, December 2023. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375823000319.

Zhang:2022:FBA

[ZWL22] Rui Zhang, Dehui Wang, and Cong Li. Flexible binomial AR(1) processes using copulas. Journal of Statistical Planning and Inference, 219(??):306–332, July 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375822000027.

Zhang:2020:MAM

[ZWYX20] Jie Zhang, Dehui Wang, Kai Yang, and Yanju Xu. A multinomial autoregressive model for finite-range time series of counts. Journal of Statistical Planning and Inference, 207(??):320–343, July 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375820300033.

Zhou:2023:CUP

[ZXS23] Yishan Zhou, Qian Xiao, and Fasheng Sun. Construction of uniform projection designs via level permutation and expansion. Journal of Statistical Planning and Inference, 222(??):209–225, January 2023. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000635.

Zhao:2022:OAB

[ZY22] Honghe Zhao and Shu Yang. Outcome-adjusted balance measure for generalized propensity score model selection. Journal of Statistical Planning and Inference, 221(??):188–200, December 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000349.

Zhang:2022:CQB

[ZYG22] Ruoyang Zhang, Yisha Yao, and Malay Ghosh. Contraction of a quasi-Bayesian model with shrinkage priors in precision matrix estimation. Journal of Statistical Planning and Inference, 221(??):154–171, December 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375822000301.

Zhou:2020:OML

[ZYL20] Weiping Zhou, Jian-Feng Yang, and Min-Qian Liu. Optimal maximin L_{2}-distance Latin hypercube designs. Journal of Statistical Planning and Inference, 207(??):113–122, July 2020. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375819301156.

Zhang:2022:GLQ

[ZYLY22] Feipeng Zhang, Jiejing Yang, Lei Liu, and Yuan Yu. Generalized linear-quadratic model with a change point due to a covariate threshold. Journal of Statistical Planning and Inference, 216(??): 194–206, January 2022. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837582100063X.

Zhao:2020:CMA

[ZZ20] Qianqian Zhao and Shengli Zhao. Constructing minimum aberration split-plot designs via complementary sets when the whole plot factors are important. Journal of Statistical Planning and Inference, 209(??):123–143, December 2020. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820300343.

Zhou:2021:IBE

[ZZ21] Niwen Zhou and Lixing Zhu. On IPW-based estimation of conditional average treatment effects. Journal of Statistical Planning and Inference, 215(??):1–22, December 2021. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375821000197.

Zhu:2024:MMA

[ZZ24] Hengkun Zhu and Guohua Zou. Mallows model averaging based on kernel regression imputation with responses missing at random. Journal of Statistical Planning and Inference, 231(??):Article 106130, July 2024. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837582300099X.

Zhu:2024:GTH

[ZZC24] Tianming Zhu, Jin-Ting Zhang, and Ming-Yen Cheng. A global test for heteroscedastic one-way FMANOVA with applications. Journal of Statistical Planning and Inference, 231(??):Article 106133, July 2024. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375823001027.

Zhang:2021:TSB

[ZZGZ21] Jin-Ting Zhang, Bu Zhou, Jia Guo, and Tianming Zhu. Twosample Behrens–Fisher problems for high-dimensional data: a normal reference approach. Journal of Statistical Planning and Inference, 213(??):142–161, July 2021. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375820301233.

Zhu:2022:FMA

[ZZWZ22] Rong Zhu, Xinyu Zhang, Alan T. K. Wan, and Guohua Zou. Frequentist model averaging under inequality constraints. Journal of Statistical Planning and Inference, 221(??):100–113, December 2022. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375822000295.